### v3

# Unstable entropy for Anosov diffeomorphisms on the $3$-torus

Abstract.
For Anosov diffeomorphisms on the $3$-torus $\mathbb{T}^{3}$ which are strongly partially hyperbolic with expanding center, we construct systems of strong unstable and center stable Margulis measures which are holonomy-invariant. This allows us to obtain a characterization of the measures of maximal unstable entropy in terms of their conditional measures along strong unstable leaves. Moreover, we show that the Margulis systems identify with Pollicott-Ruelle (co)-resonant states for the action of the diffeomorphism on $2$-forms. This shows that the unstable topological entropy and a measure of maximal unstable entropy can be retrieved from the spectral approach.
1. Introduction
1.1. Anosov diffeomorphisms with a partially hyperbolic splitting
Let $f\in\mathrm{Diffeo}^{\infty}(\mathbb{T}^{3})$ be a smooth diffeomorphism of $\mathbb{T}^{3}$. Suppose that $f$ is Anosov with splitting $T(\mathbb{T}^{3})=E_{s}\oplus E_{cu}$ where $\mathrm{dim}(E_{s})=1$ and $\mathrm{dim}(E_{cu})=2$. Suppose in addition that the unstable bundle $E_{cu}$ of $f$ splits as $E_{cu}=E_{c}\oplus E_{u}$ where $E_{c}$ is dominated by $E_{u}$, see §2 for the precise definitions. In other words, $f$ is a (strongly) partially hyperbolic diffeomorphism with splitting $T(\mathbb{T}^{3})=E_{u}\oplus E_{c}\oplus E_{s}$. We will call $E_{c}$ the center direction and $E_{u}$ the strong unstable direction of $f$. We further suppose that $f$ is transitive and orientation preserving [^1] This assumption is made to simplify the exposition. If $\mathrm{det}(f)<0$, one can consider $f\circ f$ instead. . Let $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ be the set of such diffeomorphisms.
1.2. Unstable entropy
For a partially hyperbolic diffeomorphism $f$ and an invariant probability measure $\mu$, Hu, Hua and Wu [undefw] defined the notion of metric entropy of $\mu$ with respect to the strong unstable foliation $E_{u}$, which we denote by $h^{u}_{\mu}(f)$. Another important invariant associated to $E_{u}$ is the unstable topological entropy $h^{u}_{\mathrm{top}}(f)$; an invariant introduced by Saghin and Xia [undefak] which captures the exponential growth rate of the volume of balls in the strong unstable manifold. The two notions of entropies are linked by a variational principle [undefw, Theorem D],
(1.1) $h^{u}_{\mathrm{top}}(f)=\sup_{\mu\in\mathcal{P}_{f}(\mathbb{T}^{3})}h^{u}_{\mu}(f),$
where $\mathcal{P}_{f}(\mathbb{T}^{3})$ denotes the set of $f$-invariant probability measure on $\mathbb{T}^{3}$, see §2.2 for a quick review of these notions. A probability measure which attains the supremum is called a measure of maximal $u$-entropy or $u$-MME. Their set will be denoted by $\mathcal{M}^{u}(f)$. Hu,Wu and Zhu [undefx, Theorem B] showed that $\mathcal{M}^{u}(f)$ is always non-empty, convex, compact and that its extreme points are precisely the ergodic measures of maximal $u$-entropy. In this paper, we compute $h^{u}_{\mathrm{top}}(f)$ and study some properties of $\mathcal{M}^{u}(f)$. To state our results, let us define the center Jacobian,
$J^{c}_{f}\in C^{\alpha}(\mathbb{T}^{3},\mathbb{R}),\ \forall x\in\mathbb{T}^{3},\quad J_{f}^{c}(x):=-\ln\|df_{x}|_{E_{c}(x)}\|.$
Since $J^{c}_{f}\in C^{\alpha}(\mathbb{T}^{3})$ and $f$ is Anosov there exists a unique associated equilibrium state $\mu_{J^{c}_{f}}$. Let $P(J^{c}_{f})$ be the pressure of $J^{c}_{f}$. Recall that $f$ is conjugated to its action on homology $A\in\mathrm{GL}_{3}(\mathbb{Z})$ via a Hölder continuous homeomorphism. Since, $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, the spectrum of $A$ is given by $\mathrm{Spec}(A):=\{\lambda_{s}<1<\lambda_{c}<\lambda_{u}\}$. With these notations, our first result reads:
Theorem 1 (Unstable entropy).
For any $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, one has
(1.2) $e^{h^{u}_{\mathrm{top}}(f)}=\lambda_{u}=e^{P(J^{c}_{f})},\quad\mu_{J^{c}_{f}}\in\mathcal{M}^{u}(f).$
As a consequence, $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})\ni f\mapsto h^{u}_{\mathrm{top}}(f)$ is invariant by topological conjugacy and thus locally constant.
In particular, as an equilibrium state for an Anosov map, $\mu_{J^{c}_{f}}$ satisfies exponential mixing. Moreover, $\mu_{J^{c}_{f}}$ can be expressed via Bowen’s formula in terms of periodic points, see for instance [undeft, Theorem 20.3.7].
In a recent survey [undefam], Tahzibi asked whether for an Anosov diffeomorphism which is partially hyperbolic with expanding center, a measure of maximal $u$-entropy is necessarily a measure of maximal entropy. We will say that $E_{s}$ and $E_{u}$ are jointly integrable if there is a 2 dimensional invariant foliation tangent to $E_{s}\oplus E_{u}$. Using a criteria due to Gan and Shi [undefp], we show the following result.
Corollary 1.1.
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. Then $\mu_{J^{c}_{f}}$ is a measure of maximal entropy if and only if $E_{u}$ and $E_{s}$ are jointly integrable.
For $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, joint integrability of $E_{u}$ and $E_{s}$ is equivalent to $f$ being not accessible, see [undefaq]. Accessibilty is $C^{r}$-dense ($r\geq 1$) [undefm, undeff] and $C^{1}$-open [undefl] in $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. We thus answer negatively to [undefam, Question 2] for an open and dense set of diffeomorphisms $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$.
1.3. Margulis system of measures
Next, we study more precisely the structure of $\mathcal{M}^{u}(f)$. To this end, we will construct systems of measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ (resp. $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$) on the strong unstable (resp. center stable) manifolds. The study of measures of maximal entropy (and more generally equilibrium states) via system of leaf measures has a rich and long history, see the introductions of [undefk, undefz] for an account of the existing literature.
A system of $u$-measures ((resp. cscs-measures)) is a family of Borelian measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ (resp. $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}})$ on strong unstable manifolds $(\mathcal{W}^{u}(x))_{x\in\mathbb{T}^{3}}$ (resp. center stable manifolds $(\mathcal{W}^{cs}(x))_{x\in\mathbb{T}^{3}}$), see Definition 3.3. For $p,q\in\mathbb{T}^{3}$, close enough such that $p\in\mathcal{W}^{u}(q)$ (resp. $p\in\mathcal{W}^{cs}(q)$), we denote by $\mathrm{Hol}^{u}_{p,q}:\mathcal{W}^{cs}(p)\to\mathcal{W}^{cs}(q)$ (resp. $\mathrm{Hol}^{cs}_{p,q}:\mathcal{W}^{u}(p)\to\mathcal{W}^{u}(q)$) the local $u$-holonomy (resp. $cs$-holonomy), see §2.1.4.
Theorem 2 (Margulis system).
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. There exists a system of $u$-measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ and a system of $cs$-measures $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$ such that:
  • the systems are conformal with respect to $f$
    (1.3) $\forall x\in\mathbb{T}^{3},\quad f^{*}\mu^{u}_{x}=e^{h^{u}_{\mathrm{top}}(f)}\mu^{u}_{f^{-1}(x)},\quad f_{*}\mu^{cs}_{x}=e^{h^{u}_{\mathrm{top}}(f)}\mu^{cs}_{f(x)};$
  • the systems are holonomy invariant:
    (1.4) $(\mathrm{Hol}^{cs}_{x,z})_{*}\mu^{u}_{z}=\mu^{u}_{x},\quad(\mathrm{Hol}^{u}_{y,w})_{*}\mu^{cs}_{w}=\mu^{cs}_{y},$
    for any points $x,z$ ((resp. $y,w)$ for which the center stable ((resp. strong unstable)) holonomy is well defined.
The systems are called $u$ and $cs$-Margulis systems of measures respectively. Moreover, the measure $\mu_{J^{c}_{f}}$ has a strong unstable/center stable local product structure with $u$ ((resp. $cs)$ conditionals given by $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ ((resp. $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}})$, see (5.11).
Theorem 2 should be compared to the construction of the measure of maximal entropy via systems of leaf measures. In the hyperbolic case, this is due to Sinai for maps [undefal] and to Margulis for flows [undefad]. Recently, different works have extended the construction of leaf measures to some partially hyperbolic settings, see §1.5 for a comparison with existing works. The Margulis system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ can be used to characterize the measures of maximal unstable entropy.
Corollary 1.2.
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ and $\nu\in\mathcal{P}_{f}(\mathbb{T}^{3})$. Then $\nu\in\mathcal{M}^{u}(f)$ if and only if the conditionals of $\nu$ along $\mathcal{W}^{u}$ are given by $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ ((up to a constant rescaling)), $\nu$-a.e.
The author does not know if $\mu_{J^{c}_{f}}$ is the unique measure of maximal $u$-entropy or not. We note that because of the center direction being expanding, the usual Hopf argument does not seem to give the uniqueness as in [undefk, undefh, undefan].
1.4. Pollicott-Ruelle resonances
The initial motivation of this work was to study the Pollicott-Ruelle spectrum of $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. For $0\leq k\leq 3$, let $\mathcal{D}^{\prime}_{k}(\mathbb{T}^{3})$ denote the space of $k$-currents, see §4.1 for a definition. Let us define
(1.5) $T^{*}(\mathbb{T}^{3})=E_{s}^{*}\oplus E_{cu}^{*},\quad E_{s}^{*}(E_{s})=0,\ E_{cu}^{*}(E_{cu})=0.$
One can associate to an Anosov map $f$ a discrete spectrum $\mathrm{Res}_{k}(f)\subset\mathbb{C}$ : the Pollicott-Ruelle resonances. Recall that $\lambda\in\mathrm{Res}_{k}(f)$ if and only if
(1.6) $\exists u\in\mathcal{D}^{\prime}_{k}(\mathbb{T}^{3})\setminus\{0\},\quad f_{*}u=\lambda u,\ \mathrm{WF}(u)\subset E_{cu}^{*}.$
Here, $\mathrm{WF}(u)\subset T^{*}\mathbb{T}^{3}\setminus\{0\}$ denotes the wavefront set of $u$, see [undefv, Chapter VIII]. In this case, the current $u$ is called a resonant state associated to the resonance $\lambda$. We have a dual notion of co-resonant state:
(1.7) $\exists u\in\mathcal{D}^{\prime}_{3-k}(\mathbb{T}^{3})\setminus\{0\},\quad f^{*}u=\lambda u,\ \mathrm{WF}(u)\subset E_{s}^{*}.$
We say that $\lambda\in\mathbb{R}_{+}$ is the first resonance for the action on $k$-forms if $\lambda\in\mathrm{Res}_{k}(f)$ and $\mathrm{Res}_{k}(f)\subset\{z\in\mathbb{C},|z|\leq\lambda\}.$ For $k=0$ (resp. $k=1$), it is known that the first resonance is given by $1$ (resp. $e^{h_{\mathrm{top}}(f)}$) and that the associated (co)-resonant states can be used to reconstruct the SRB measure (resp. measure of maximal entropy), see [undefa, undefq] for Anosov maps and [undefg, undefz] for Anosov flows. For $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, it is therefore natural to ask what is the first resonance for the action on $2$-forms and whether the corresponding (co)-resonant states have dynamical meaning or not.
Theorem 3 (First Pollicott-Ruelle resonance).
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. Then $e^{h^{u}_{\mathrm{top}}(f)}$ is the first Pollicott-Ruelle resonance for the action on $2$-forms. Moreover, $e^{h^{u}_{\mathrm{top}}(f)}$ is the only resonance on the critical circle $\{z\in\mathbb{C},|z|=e^{h^{u}_{\mathrm{top}}(f)}\}$, it does not have Jordan block and the space of resonant ((resp. co-resonant)) states is one-dimensional.
Let $\theta$ ((resp. $\nu)$ be a corresponding resonant ((resp. co-resonant)) state associated to $e^{h^{u}_{\mathrm{top}}(f)}$. Then one has
(1.8) $\forall x\in\mathbb{T}^{3},\quad\theta_{|\mathcal{W}^{cs}(x)}=\mu^{cs}_{x},\ \nu_{|\mathcal{W}^{u}(x)}=\mu^{u}_{x},$
where $g|_{\mathcal{N}}$ denotes the restriction of a current $g$ to a submanifold $\mathcal{N}.$222see for instance [undefv, Corollary 8.2.7]. Finally, the trace of the spectral projector at the resonance $\lambda=e^{h^{u}_{\mathrm{top}}(f)}$ is given by $\mu_{J^{c}_{f}}$.
In other words, one can recover the unstable topological entropy as a Pollicott-Ruelle resonance. The (co)-resonant states identify with Margulis systems of measures and $\mu_{J^{c}_{f}}$ can be reconstructed as the trace of the spectral projector at the first resonance.
Remark 1.
The restrictions in (1.8) are well defined by wavefront set conditions. In the case of the co-resonant state, we show in fact the stronger inclusion $\mathrm{WF}(\nu)\subset E_{cs}^{*}\subset E_{s}^{*}$, see (5.13), since the condition in (1.7) is a priori not enough to restrict a co-resonant state to a strong unstable manifold $\mathcal{W}^{u}(x)$.
1.5. Comparison with existing results.
1.5.1. Unstable entropy
The topological unstable entropy was shown to be upper semi-continuous by Yang [undefap] and continuous in our setting by Wu [undefao]. That $h^{u}_{\mathrm{top}}(f)$ is constant in a neighborhood of a linear automorphism $A$ follows from the work of Hua, Saghin and Xia [undefy]. Here, we extend this observation to all Anosov maps $f$ isotopic to $A$. The facts that the unstable topological entropy coincides with the pressure of $J^{c}_{f}$ and that $\mu_{J^{c}_{f}}\in\mathcal{M}^{u}(f)$ seem to be new, to the best of the author’s knowledge.
1.5.2. Systems of Margulis measures
Recently, different works have constructed systems of leaf measures for some classes of partially hyperbolic systems in order to study measures of maximal entropy, and more generally equilibrium states. Most relevant to our paper are the constructions of Climenhaga, Pesin and Zelerowicz [undefk], Parmenter and Pollicott [undefae], Carrasco and Rodriguez-Hertz [undefi], Buzzi, Fisher and Tazhibi [undefh] and Ures, Viana, Yang and Yang [undefan].
We emphasize that Theorem 2 does not follow from any of the previously cited works. Indeed, for $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, since the center direction $E_{c}$ is uniformly expanded, $f$ does not satisfy the Lyapunov stability condition of [undefk, undefae], it is not a center isometry as required in [undefi], nor it is a flow type diffeomorphism as in [undefh]. Although any $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is topologically conjugated to an Anosov automorphism $A$ of $\mathbb{T}^{3}$, it does not necessarily factor over Anosov as in [undefan]. Indeed, as shown in [undefaq], the conjugacy maps the strong unstable foliation of $f$ to that of $A$ if and only if $E_{u}$ and $E_{s}$ are jointly integrable. As stated in the introduction, this condition fails for a dense and open subset of $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. For such $f$, the Margulis system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ cannot be obtained by simply pulling back by the conjugacy the Lebesgue measure of $\mathcal{W}^{u}_{A}$, where $\mathcal{W}^{u}_{A}$ is the strong unstable manifold of $A$. When $f$ factor over Anosov, Corollary 1.2 does follow from [undefan, Theorem B] but since $f$ does not have c-mostly contracting center, [undefan, Theorem C] does not apply and the uniqueness of the $u$-MME is open even in this case. For a linear automorphism $A\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, the foliations are linear and are in particular absolutely continuous. This means that using Corollary 1.2 and the standard Hopf argument, one deduces that $A$ has a unique $u$-MME, see for instance [undefh, Proposition 3.8].
To the best of the author’s knowledge, Theorem 2 is the first general construction of Margulis systems for a class of partially hyperbolic systems with uniformly expanding center. From a technical point of view, this makes showing the $\mathrm{Hol}^{cs}$-invariance (1.4) of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ much harder due to the lack of regularity of center stable holonomies. To overcome this difficulty, we combine dynamical ideas inspired by [undefh, undefi, undefk] with functional techniques as in [undefz, undefaa], see the next subsection for an outline of the argument. We note that Alvarez, Leguil, Obata and Santiago studied $u$-Gibbs measures for $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ [^3] Their techniques actually allow them to study $C^{2}$ diffeomorphisms rather than smooth ones. in [undef]. The lack of regularity of center stable holonomies is also aknowledged as a source of technical difficulty in their work, see [undef, §2.2.3], although they use different techniques to overcome it.
1.5.3. Pollicott-Ruelle resonances
Theorem 3 seems to be the first result on the first resonance for Anosov maps for $k\neq 0,d_{s}$ where $d_{s}=\mathrm{dim}(E_{s})$. The close link between (co)-resonant states for $k=0,d_{s}$ and systems of leaf measures for Anosov diffeomorphism is well understood [undefq]. This was used by the author in [undefz] to reprove classical facts on the thermodynamical formalism for Anosov flows.
Recently, the author combined the functional approach for general Anosov actions obtained by Guedes Bonthonneau, Guillarmou, Hilgert, and Weich [undefr, undefs] and the work of Carrasco and Rodriguez-Hertz [undefi] to study the measure of maximal entropy for minimal Anosov actions [undefaa]. The present work can be seen as another instance of partially hyperbolic systems for which certain important dynamical invariant and measure can be reconstructed from the spectral approach.
We note that the functional approach was also used to construct physical and SRB measures for some partially hyperbolic systems, see [undefj] and the references therein.
1.6. Structure of the paper
In §2, we define the class $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ and recall some facts about the dynamical foliations tangent to the invariant bundles $E_{\bullet}$, for $\bullet=s,c,u,cs,cu$ and the local product structure they define. We recall as well the definition of the conditional entropy of an invariant measure along the strong unstable foliation and prove a basic lemma about the pressure of the center Jacobian.
In §3, we construct the unstable system of Margulis measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ using a compactness argument inspired by [undefh, undefi]. The compactness used in the argument follows from a precise volume bound (3.4) which is a consequence of the work of Potrie [undefag]. We emphasize that since the complementary foliation $\mathcal{W}^{cs}$ is not contracting, this construction does not suffice to obtain the holonomy invariance (1.4) of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ as in [undefh, undefi]. Nevertheless, using a Carathéodory construction inspired by Climenhaga, Pesin and Zelerowicz [undefk], we show that $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ is measurable using which allows us to deduce Corollary 1.2 using an argument of Tahzibi [undefam].
In §4, we recall some facts about Pollicott-Ruelle resonances and show Theorem 3, except for (1.8), using the functional approach. The argument is shorter than in [undefz, undefaa] since $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is conjugated to a linear model, which we can use to simplify some proofs. We also deduce Theorem 1 and Corollary 1.1.
Finally, §5 is devoted to the construction of the $cs$-system and the proofs of (1.4) and (1.8). We use a combination of functional and dynamical ideas. In summary:
  • the $cs$-system $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$ is constructed from the resonant state using the functional approach, see Proposition 5.1. This establishes (1.4) and (1.8) for $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$;
  • we show using the functional approach that $\mu_{J^{c}_{f}}$ can be obtained as the limit of pullbacks of $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$, see Proposition 5.3. This should be compared with [undefk, Theorem 4.7 (1)] and [undefae, Theorem 1.1] although our techniques are quite different;
  • we deduce that $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$ is equal to the $cs$-conditionals of $\mu_{J^{c}_{f}}$ mimicking the dynamical argument of [undefk, Lemma 8.1];
  • using an adaptation of [undefk, Lemma 2.12], the local product structure of $\mu_{J^{c}_{f}}$ and Corollary 1.2, we deduce the holonomy invariance (1.4) of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ from the holonomy invariance of $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$. This last argument seems to be the main novelty compared to previous works and might be of independent interest;
  • finally, we show (1.8) for $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ by adapting [undefz, Lemma 3.2].
Acknowledgements. The author would like to first thank Colin Guillarmou and Thibault Lefeuvre for their guidance and advice during the writing of this paper. The author would also like to thank Jérôme Buzzi, Sylvain Crovisier, François LeDrappier, Ali Tahzibi and Amie Wilkinson for answering some questions related to the project.
The author would like to further thank Ali Tahzibi for noticing that the previous argument given in Remark 3 was incomplete and Jérôme Buzzi for making him aware of a gap in the previous proof of the measurability of the Margulis system in Proposition 3.4. These issues have now been resolved in the current version of the paper. Finally, he would like to thank Ali Tahzibi for pointing out to him that the $u$-MME is unique for linear automorphisms and Gabriel Rivière for pointing out the work [undefaj], which lead to Remark 2.
This research was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement no. 101162990 — ADG).
2. Preliminaries
2.1. Anosov diffeomorphism with a partially hyperbolic splitting
2.1.1. Definition
Let $f\in\mathrm{Diffeo}^{\infty}(\mathbb{T}^{3})$ be a smooth diffeomorphism. We say that $f$ is Anosov and strongly partially hyperbolic with expanding center if there exists a continuous splitting of the tangent bundle $T(\mathbb{T}^{3})=E_{u}\oplus E_{c}\oplus E_{s},$ which is $df$-invariant and a Riemannian metric $\|.\|$ adapted to the splitting such that
$x\mapsto\lambda_{x}^{\bullet}:=\|df(x)|_{E_{\bullet}}\|,\quad\bullet=s,c,u,$
are continuous and satisfy $\lambda_{x}^{s}<1<\lambda_{x}^{c}<\lambda_{x}^{u}$ for any $x\in\mathbb{T}^{3}$. In particular, $f$ is an Anosov map with stable bundle $E_{s}$ and unstable bundle $E_{cu}:=E_{c}\oplus E_{u}$. The bundles $E_{c}$ and $E_{u}$ are respectively called the center and strong unstable bundle of $f$. Finally, $E_{cs}:=E_{c}\oplus E_{s}$ is the center stable bundle of $f$. For a detailed introduction to Anosov diffeomorphisms, we refer to [undeft].
We will denote by $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ the set of such diffeomorphisms which furthermore preserve the orientation of $\mathbb{T}^{3}$. It is easily seen to be $C^{1}$-open and examples are provided by $A\in\mathrm{GL}_{3}(\mathbb{Z})$ with real distincts eigenvalues $\lambda_{s}<1<\lambda_{c}<\lambda_{u}$.
2.1.2. Invariant manifolds
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. Since $f$ is Anosov, $E_{s}$ and $E_{cu}$ are integrable to $f$-invariant foliations $\mathcal{W}^{s}$ and $\mathcal{W}^{cu}$ called respectively the stable and unstable foliations of $f$, see [undeft, Theorem 6.2.3]. Moreover, since $E_{u}$ is the strong unstable foliation of a strongly partially hyperbolic diffeomorphism, it integrates to a $f$-invariant foliation $\mathcal{W}^{u}$ called the strong unstable foliation of $f$, see [undefaf, Theorem 4.1]. Actually, by the work of Brin, Burago and Ivanov [undefe] and Potrie [undefag], any $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is dynamically coherent. This means that the center stable bundle $E_{cs}$ also integrates to a $f$-invariant foliation $\mathcal{W}^{cs}$ called the center stable foliation of $f$. The collection of leaves $\mathcal{W}^{c}(x):=\mathcal{W}^{cs}(x)\cap\mathcal{W}^{cu}(x)$ for $x\in\mathbb{T}^{3}$ defines the center foliation which integrates the center bundle $E_{c}$. For any $\bullet=s,c,u,cs,cu$, any $x\in\mathbb{T}^{3}$ and any $\delta>0$, let $\mathcal{W}^{\bullet}(x,\delta):=\{y\in\mathcal{W}^{\bullet}(x)\mid d^{\bullet}(x,y)<\delta\}$, where $d^{\bullet}$ is the distance on the leaf.
2.1.3. Local product structure and rectangles
By transversality of the foliations $\mathcal{W}^{cs}$ and $\mathcal{W}^{u}$, there is an $\varepsilon_{0}>0$ and a $C>0$ such that
(2.1) $\forall x,y\in\mathbb{T}^{3},\ d(x,y)<\varepsilon<\varepsilon_{0}\ \Rightarrow\ \mathcal{W}^{cs}_{C\varepsilon}(x)\cap\mathcal{W}^{u}_{C\varepsilon}(y)=\{[x,y]\},$
where $[x,y]$ is the Bowen bracket of $x$ and $y$.
Definition 2.1.
A closed set $R\subset\mathbb{T}^{3}$ is called a rectangle if for every $x,y\in R$, the Bowen bracket $[x,y]$ exists and $x,y]\in R$.
Rectangles exist because of the local product structure, see [undefk, Eq. $(2.2)$].
2.1.4. Holonomies
We now define the center stable and unstable holonomies. Let $R$ be a rectangle and $x\in R$. We will denote by $\mathcal{W}^{\bullet}_{R}(x),$ for $\bullet=cs,u,$ the connected component of $\mathcal{W}^{\bullet}(x)\cap R$ containing $x$. For $x,y\in R$, the strong unstable holonomy is
(2.2) $\mathrm{Hol}^{u}_{x,y}:\mathcal{W}^{cs}_{R}(x)\to\mathcal{W}^{cs}_{R}(y),\quad\mathrm{Hol}^{u}_{x,y}(z):=[y,z]\in\mathcal{W}^{cs}(y)\cap\mathcal{W}^{u}(z).$
The center stable holonomy is
(2.3) $\mathrm{Hol}^{cs}_{x,y}:\mathcal{W}^{u}_{R}(x)\to\mathcal{W}^{u}_{R}(y),\quad\mathrm{Hol}^{cs}_{x,y}(z):=[z,y]\in\mathcal{W}^{cs}(z)\cap\mathcal{W}^{u}(y).$
Given $x,y\in R$ and $A\subset\mathcal{W}^{cs}_{R}(x),B\subset\mathcal{W}^{cs}_{R}(y)$, we say that $A$ and $B$ are $\mathrm{Hol}^{u}$-equivalent if $\mathrm{Hol}^{u}_{x,y}(A)=B$. Given $\delta>0$, we say that $A$ and $B$ are $\delta$-equivalent if the size of the holonomy is bounded by $\delta$. We define likewise the notion of $\mathrm{Hol}^{cs}$-equivalence. Two functions $\psi\in C^{\infty}(\mathcal{W}^{cs}_{R}(x))$ and $\phi\in C^{\infty}(\mathcal{W}^{cs}_{R}(y))$ are said to be $\mathrm{Hol}^{u}$-equivalent (resp. $\delta$-equivalent) if their supports are $\mathrm{Hol}^{u}$-equivalent (resp. $\delta$-equivalent) and $\phi\circ\mathrm{Hol}^{u}_{x,y}=\psi$ on the support of $\psi$. We define likewise the notions of $\mathrm{Hol}^{cs}$-equivalence (resp. $\delta$-equivalence) for functions on strong unstable leaves.
The strong unstable foliation is absolutely continuous, see (5.5) for the exact statement we will need. This fact will be used in the proof of Proposition 5.2 to show the $\mathrm{Hol}^{u}$-invariance of $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$. The center stable foliation is not always absolutely continuous and this explains why proving the $\mathrm{Hol}^{cs}$-invariance of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ is harder. We refer to [undef, §2.2.3] for a more detailed discussion of the lack of regularity of center stable holonomies.
2.2. Unstable entropy
2.2.1. Conditional measures
We recall some facts about measurable partitions and conditional measures and refer to [undefu] for more details. Let $(X,\mu)$ be a probability space. For $\xi$ a partition of $X$ into $\mu$-measurable sets, we write $\xi(x)\in\xi$ for the element of the partition containing $x$. Let $\xi_{1},\xi_{2}$ be two partitions. We say that $\xi_{1}$ refines $\xi_{2}$ denoted $\xi_{2}\prec\xi_{1}$ if for $\mu$-a.e. $x\in X$, one has $\xi_{1}(x)\subset\xi_{2}(x)$. The joint of $\xi_{1}$ and $\xi_{2}$ is the partition defined by $\xi_{1}\vee\xi_{2}:=\{\xi_{1}(x)\cap\xi_{2}(x)\mid x\in X\}$. It is the least fine partition that refines both $\xi_{1}$ and $\xi_{2}$. A partition $\xi$ is said to be measurable if there is a sequence of finite partitions $\xi_{n}$ such that $\xi=\vee_{n=0}^{+\infty}\xi_{n}.$
For a measurable partition $\xi$, Rokhlin showed that there exists a system of conditional measures $(\mu_{x}^{\xi})_{x\in X}$ such that:
  • for any $x\in X$, $\mu_{x}^{\xi}$ is a probability measure on $\xi(x)$,
  • if $\xi(x)=\xi(y)$, then $\mu_{x}^{\xi}=\mu_{y}^{\xi}$,
  • for any $\psi\in L^{1}(X,\mu)$, one has
    (2.4) $\int_{X}\psi d\mu=\int_{X}\int_{\xi(x)}\psi(z)d\mu^{\xi}_{x}(z)d\mu(x).$
Moreover, the system of conditional measures is unique $\mu$-mod 0, that is, two systems of conditional measures coincide for $\mu$-a.e. $x\in X$.
The main example we will consider is $X=R$ a rectangle and $(\xi(x))_{x\in R}=(\mathcal{W}^{cs}_{R}(x))_{x\in R}$ the partition into center stable manifolds. This partition is easily seen to be measurable. For this partition, we will write $\mu^{cs}_{x}=\mu_{x}^{\xi}$ and refer to $(\mu^{cs}_{x})_{x\in R}$ as the conditionals of $\mu$ on the center stable manifolds. For $q\in R$, define a measure $\tilde{\mu}_{q}$ on $\mathcal{W}^{u}_{R}(q)$ by
(2.5) $\forall E\subset\mathcal{W}^{u}_{R}(q),\quad\tilde{\mu}_{q}(E):=\mu(\cup_{y\in E}\mathcal{W}^{cs}_{R}(y)).$
Then (2.4) can be rewritten for any $q\in R$ as,
(2.6) $\forall\phi\in L^{1}(R,\mu),\quad\mu(\phi)=\int_{\mathcal{W}^{u}_{R}(q)}\int_{\mathcal{W}^{cs}_{R}(y)}\phi(z)d\mu^{cs}_{y}(z)d\tilde{\mu}^{u}_{q}(y),$
see [undefk, Equation $(2.6)$]. One defines the conditionals of $\mu$ on the strong unstable manifolds $(\mu^{u}_{x})_{x\in R}$ by replacing $cs$ by $u$ in the previous definition. Although this is not explicit in the notations, the conditionals depend on the rectangle $R$. Note however that the ambiguity only consists in a normalizing constant, see [undefk, Lemma 2.10].
2.2.2. Metric unstable entropy
We now recall the definition of the metric unstable entropy. For further details, we refer to [undefw]. In this subsection, we consider $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ and $\mu\in\mathcal{P}_{f}(\mathbb{T}^{3})$ a $f$-invariant probability measure. For a partition $\xi$, define its diameter to be $\mathrm{diam}(\xi):=\sup_{x\in\mathbb{T}^{3}}\mathrm{diam}(\xi(x))$. Fix a small $\varepsilon_{0}>0$ and let $\mathcal{Q}$ be the set of $\mu$-measurable partitions of diameter less than $\varepsilon_{0}$. For $\xi\in\mathcal{Q}$, we define a finer partition $\eta=:\mathcal{Q}^{u}(\xi)$ subordinated to the unstable foliation by defining $\eta(x)$ to be the connected component of $\xi(x)\cap\mathcal{W}^{u}(x)$ containing $x$, for any $x\in\mathbb{T}^{3}$. Let $\mathcal{Q}^{u}$ denote the set of all partitions obtained this way. The conditional entropy of $\xi$ given $\eta$ with respect to $\mu$ is given by
$H_{\mu}(\xi\mid\eta):=-\int_{\mathbb{T}^{3}}\ln(\mu^{\eta}_{x}(\xi(x)))d\mu(x).$
For $m,n\in\mathbb{Z}$, define $\xi_{m}^{n}:=\vee_{i=m}^{n}f^{-i}\xi$. The conditional entropy of $f$ with respect to a measurable partition $\xi$ given $\eta$ is
$h_{\mu}(f,\xi\mid\eta)=\limsup_{n\to+\infty}\frac{1}{n}H_{\mu}(\xi_{0}^{n}\mid\eta).$
Suppose furthermore that $\mu$ is ergodic, then the unstable metric entropy of $f$ with respect to $\mu$ is defined to be
(2.7) $\forall\xi\in\mathcal{Q},\ \forall\eta\in\mathcal{Q}^{u},\quad h^{u}_{\mu}(f):=h_{\mu}(f,\xi\mid\eta),$
where the value does not depend on $\xi$ or $\eta$, see [undefw, Theorem A].
2.2.3. Topological unstable entropy and variational principle.
The topological counter part of the metric unstable entropy is the topological unstable entropy. It was introduced by Saghin and Xia [undefak]. It is equal to the exponential growth rate of the volume of strong unstable balls:
(2.8) $\forall x\in\mathbb{T}^{3},\ \forall\delta>0,\quad h^{u}_{\mathrm{top}}(f):=\limsup_{n\to+\infty}\frac{1}{n}\ln\mathrm{Vol}\big(f^{n}(\mathcal{W}^{u}(x,\delta))\big),$
where $\mathrm{Vol}$ is the induced volume by the Riemannian metric. The limit does not depend on either, $x\in\mathbb{T}^{3}$, $\delta>0$ or the choice of the Riemannian metric. Hu, Hua and Wu [undefw, Definition 1.4] gave an alternative definition of $h^{u}_{\mathrm{top}}(f)$ in terms of spanning or separated sets which we will not need in this paper. The link between the two notions of unstable entropies is given by the variational principle (1.1), see [undefw, Theorem D]. A measure $\mu\in\mathcal{P}_{f}(\mathbb{T}^{3})$ is a measure of maximal unstable entropy (or $u$-MME) if $h^{u}_{\mathrm{top}}(f)=h^{u}_{\mu}(f)$. Their set is denoted by $\mathcal{M}^{u}(f)$. By the work of Hu, Wu and Zhu [undefx, Theorem B], the set $\mathcal{M}^{u}(f)$ is always non-empty, convex and compact. Moreover, its extreme points are exactly the ergodic $u$-MMEs.
2.3. Center Jacobian
In this subsection, we define the center Jacobian. Let
(2.9) $J^{c}_{f}\in C^{\alpha}(\mathbb{T}^{3},\mathbb{R}),\ \forall x\in\mathbb{T}^{3},\quad J_{f}^{c}(x):=-\ln\|df_{x}|_{E_{c}(x)}\|,$
where $\|.\|$ denotes the adapted Riemannian metric and $\alpha>0$. By the thermodynamical formalism (see [undeft, Theorem 20.3.7]), there is a unique equilibrium state $\mu_{J^{c}_{f}}$ associated to $J_{f}^{c}$. It is the unique invariant probability measure such that
(2.10) $h_{\mu_{J^{c}_{f}}}(f)+\int_{\mathbb{T}^{3}}J^{c}_{f}(p)d\mu_{J^{c}_{f}}(p)=P(J^{c}_{f}):=\sup_{\mu\in\mathcal{P}_{f}(\mathbb{T}^{3})}h_{\mu}(f)+\int_{\mathbb{T}^{3}}J^{c}_{f}(p)d\mu(p),$
where $P(J^{c}_{f})$ is the pressure of the center Jacobian and $h_{\mu}(f)$ is the metric entropy of the map $f$ with respect to $\mu$. We show an inequality between $h^{u}_{\mathrm{top}}(f)$ and $P(J^{c}_{f})$.
Lemma 2.2.
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. Then one has
(2.11) $h^{u}_{\mathrm{top}}(f)\geq P(J^{c}_{f}).$
Proof.
We use [undefw, Corollary A.1] which states that
(2.12) $h_{\mu_{J^{c}_{f}}}(f)\leq h^{u}_{\mu_{J^{c}_{f}}}(f)+\lambda_{c}(f,\mu_{J^{c}_{f}}),$
where $\lambda_{c}(f,\mu_{J^{c}_{f}})$ is the central Lyapunov exponent of $\mu_{J^{c}_{f}}$. Recall that it satisfies
$\text{for }\mu_{J^{c}_{f}}\text{-a.e }x\in\mathbb{T}^{3},\ \forall v\in E_{c}(x),\ \quad\lim_{n\to+\infty}\frac{1}{n}\ln\|df^{n}_{x}(v)\|=\lambda_{c}(f,\mu_{J^{c}_{f}}).$
Now, the measure $\mu_{J^{c}_{f}}$ is ergodic so by the ergodic theorem and the chain rule,
$\text{for }\mu_{J^{c}_{f}}\text{-a.e }x\in\mathbb{T}^{3},\quad\lim_{n\to+\infty}-\frac{1}{n}\ln\|df^{n}_{x}|_{E_{c}(x)}\|=\lim_{n\to+\infty}\frac{1}{n}\sum_{k=0}^{n-1}J^{c}_{f}(f^{k}x)=\int_{\mathbb{T}^{3}}J^{c}_{f}d\mu_{J^{c}_{f}}.$
Plugging this into (2.12) and using (2.10) yields
(2.13) $h^{u}_{\mu_{J^{c}_{f}}}(f)\geq h_{\mu_{J^{c}_{f}}}(f)-\lambda_{c}(f,\mu_{J^{c}_{f}})=h_{\mu_{J^{c}_{f}}}(f)+\int_{\mathbb{T}^{3}}J^{c}_{f}d\mu_{J^{c}_{f}}=P(J^{c}_{f}).$
Using $h^{u}_{\mu_{J^{c}_{f}}}(f)\leq h^{u}_{\mathrm{top}}(f)$ gives (2.11). ∎
3. Construction of a Margulis system of measures
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. In this section, we construct a system of $u$-measures for $f$. Recall that there exists a Hölder continuous homeomorphism $h:\mathbb{T}^{3}\to\mathbb{T}^{3}$ and a hyperbolic matrix $A\in\mathrm{GL}_{3}(\mathbb{Z})$ such that $h\circ f=A\circ h,$ see for instance [undeft, Theorem 18.6.1]. The matrix $A$ has three distinct and real eigenvalues $\lambda_{s}<1<\lambda_{c}<\lambda_{u}$ and the corresponding eigenlines are denoted by $E_{s}^{A}$, $E_{c}^{A}$ and $E_{u}^{A}$ respectively. We have $A\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ and the corresponding foliations are given by $\mathcal{W}^{\bullet}_{A}(x)=x+E_{\bullet}^{A}$, for $\bullet=s,c,u$.
Let $H$, $F$ and $A$ denote a lift to $\mathbb{R}^{3}$ of $h,f$ and $A$ respectively. For a foliation $\mathcal{N}$ of $\mathbb{T}^{3}$, we denote by $\widetilde{\mathcal{N}}$ the lifted foliation of $\mathbb{R}^{3}$. Then one has:
  • the conjugacy lifts to $\mathbb{R}^{3}$, that is,
    (3.1) $H\circ F=A\circ H;$
  • the conjugacy $H$ is isotopic to $\mathrm{Id}$. In particular, there exists $L>0$ such that
    (3.2) $\|H-\mathrm{Id}\|_{\infty}:=\sup_{x\in\mathbb{R}^{3}}d(H(x),x)\leq L,$
    where $d$ denotes the Euclidean distance on $\mathbb{R}^{3}$;
  • for any $\bullet=s,c,cs,cu$, one has
    (3.3) $);\forall\tilde{x}\in\mathbb{R}^{3},\quad H\big(\widetilde{\mathcal{W}_{f}^{\bullet}}(\tilde{x})\big)=\widetilde{\mathcal{W}^{\bullet}_{A}}(H(\tilde{x}$));
  • one has $)H\big(\widetilde{\mathcal{W}_{f}^{u}}(\tilde{x})\big)=\widetilde{\mathcal{W}^{u}_{A}}(H(\tilde{x}$)) for any $\tilde{x}\in\mathbb{R}^{3}$ if and only if $E_{s}$ and $E_{u}$ are jointly integrable.
The first three facts were proven by Potrie in [undefag] and the last one was proven by Ren, Gan and Zhang in [undefaq]. Note that the fact that $E_{s}$ and $E_{u}$ are not jointly integrable is an open and dense condition in $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, see [undefm, undeff, undefl].
We first show the following uniform volume bounds.
Proposition 3.1.
Let $\delta>0$. There exists $C_{1}(\delta),C_{2}(\delta)>0$ such that for any $\tilde{x}\in\mathbb{R}^{3}$,
(3.4) $\forall n\in\mathbb{N},\quad C_{1}(\delta)\leq\lambda_{u}^{-n}\mathrm{Vol}(F^{n}\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))\leq C_{2}(\delta).$
Proof.
By Potrie [undefag, Corollary 7.7], the strong unstable foliation is quasi-isometric:
(3.5) $\exists C>0,\ \forall\tilde{x}\in\mathbb{R}^{3},\ \forall\tilde{y},\tilde{z}\in\widetilde{W}_{f}^{u}(\tilde{x}),\quad d(\tilde{y},\tilde{z})\leq d^{u}(\tilde{y},\tilde{z})\leq Cd(\tilde{y},\tilde{z})+C,$
[…]
$\mathrm{diam}(F^{n}\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))\leq\mathrm{Vol}(F^{n}\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))\leq C\mathrm{diam}(F^{n}\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))+C,$
where the diameter is taken with respect to $d$. Now, using (3.2), we obtain
$\mathrm{diam}(H\circ F^{n}\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))-L\leq\mathrm{Vol}(F^{n}\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))\leq C\mathrm{diam}(H\circ F^{n}\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))+C(1+L).$
Using (3.1), this then implies
$\mathrm{diam}(A^{n}\circ H\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))-L\leq\mathrm{Vol}(F^{n}\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))\leq C\mathrm{diam}(A^{n}\circ H\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))+C(1+L).$
Since $\lambda_{u}$ is simple and the largest eigenvalue of $A$, there is $K>0$ such that for any $\tilde{x},\tilde{y}\in\mathbb{R}^{3},$ one has $\|A^{n}(\tilde{x}-\tilde{y})\|\leq K\lambda_{u}^{n}\|\tilde{x}-\tilde{y}\|$. In particular, using (3.2),
$C\mathrm{diam}(A^{n}\circ H\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))\leq K\lambda_{u}^{n}\mathrm{diam}(H\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))\leq K(2\delta+L)\lambda_{u}^{n}.$
[…]
(3.6) $\mathrm{diam}\big(\pi^{u}_{\tilde{x}}\big(H\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta)\big)\big)\geq c(\delta).$
Indeed, this follows from the transversality of $E_{c}$ and $E_{u}$, the continuity of the bundles and the compactness of $\mathbb{T}^{3}.$ In particular, one has
$\mathrm{diam}(A^{n}\circ H\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta))\geq\mathrm{diam}\big(\pi_{\tilde{x}}^{u}\big(H\widetilde{\mathcal{W}}_{f}^{u}(\tilde{x},\delta)\big)\big)\lambda_{u}^{n}\geq c(\delta)\lambda_{u}^{n}.$
This gives the lower bound in (3.4). ∎
Combining (2.8) and (3.4), and since the volume growth on the lifted foliation is equal to the volume growth on the strong unstable foliation, we obtain:
Corollary 3.2.
For any $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, one has
(3.7) $e^{h^{u}_{\mathrm{top}}(f)}=\lambda_{u}.$
We now construct the systems of $u$-measures. We start by a definition.
Definition 3.3 (System of measures).
For $\bullet=u,cs$, a system of $\bullet$-measures $\mu^{\bullet}$ is a family $\mu^{\bullet}=(\mu^{\bullet}_{x})_{x\in\mathbb{T}^{3}}$ such that:
  1. (1)
    for any $x\in\mathbb{T}^{3}$, $\mu^{\bullet}_{x}$ is a Borelian measure on $\mathcal{W}^{\bullet}(x)$ which is finite on compact subsets of $\mathcal{W}^{\bullet}(x)$;
  2. (2)
    for any $x,y\in\mathbb{T}^{3}$ such that $\mathcal{W}^{\bullet}(x)=\mathcal{W}^{\bullet}(y)$, one has $\mu^{\bullet}_{x}=\mu^{\bullet}_{y}$.
We say that $\mu^{\bullet}$ is a measurable ((resp. continuous)) system of measures if $\mathbb{T}^{3}$ is covered by foliation charts $B$ such that for any $\phi\in C_{c}^{\infty}(B,\mathbb{R})$, the mapping $x\mapsto\mu^{\bullet}_{x}(\phi_{|\mathcal{W}^{\bullet}_{B}}(x))$ is measurable ((resp. continuous)). Here, $\mathcal{W}^{\bullet}_{B}(x)$ denotes the connected component of $\mathcal{W}^{\bullet}(x)\cap B$ containing $x$.
We start by constructing a measurable system of $u$-measure which has the scaling property (1.3). The holonomy-invariance (1.4) will be proved in §5.
Proposition 3.4 (Measurable system of $u$-measures).
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. Then there exists a measurable system of $u$-measure $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ such that:
  • for any $x\in\mathbb{T}^{3}$, $\mu_{x}^{u}$ has full support in $\mathcal{W}^{u}(x)$;
  • one has the scaling property
    (3.8) $\forall x\in\mathbb{T}^{3},\quad f^{*}\mu^{u}_{x}=e^{h^{u}_{\mathrm{top}}(f)}\mu^{u}_{f^{-1}(x)}.$
Proof.
We use a compactness argument similar to [undefh, undefi] and first construct a system of measure which has full support in each leaf and has the scaling property.
There exists a system of $u$-measures with full support in each leaf and the scaling property (3.8). A test function $\psi$ is the data of $(\psi_{x})_{x\in\mathbb{T}^{3}}$ where for any $x\in\mathbb{T}^{3}$, $\psi_{x}\in C_{c}(\mathcal{W}^{u}(x))$ is a continuous function with compact support in $\mathcal{W}^{u}(x)$. The set of test functions is denoted by $\mathrm{Con}^{u}(f)$. Let $\mathrm{Meas}^{u}(f)$ be the set of systems of $u$-measures. We endow $\mathrm{Meas}^{u}(f)$ with the weak topology induced by $\mathrm{Con}^{u}(f)$. We define $\mathrm{Con}_{+}^{u}(f)$ to be the subset of test functions $\psi$ such that for any $x\in\mathbb{T}^{3}$, $\psi_{x}$ is non-negative and non identically zero. Denote by $\mathrm{Vol}_{\mathcal{W}^{u}}\in\mathrm{Meas}^{u}(f)$ the system of measures induced by the Lebesgue measure on each unstable manifold. Define
(3.9) $\mathcal{X}:=\overline{\mathrm{Conv}\big\{\nu^{n}:=e^{-nh^{u}_{\mathrm{top}}(f)}(f^{*})^{n}\mathrm{Vol}_{\mathcal{W}^{u}}\mid n\in\mathbb{N}\big\}}\subset\mathrm{Meas}^{u}(f),$
where $\mathrm{Conv}(X)$ denotes the convex hull of $X$ and $\overline{X}$ its closure.
Lemma 3.5.
The subset $\mathcal{X}\subset\mathrm{Meas}^{u}(f)$ is compact.
Proof.
This follows from Proposition 3.1. Indeed, we first show that given $x_{1},x_{2}\in\mathbb{T}^{3}$, $X_{1}\subset\mathcal{W}^{u}(x_{1})$ and $X_{2}\subset\mathcal{W}^{u}(x_{2})$ two open and pre-compact sets, there is a constant $\hat{e}(X_{1},X_{2})>0$ such that
(3.10) $\forall n\geq 0,\quad\frac{1}{\hat{e}(X_{1},X_{2})}\leq\frac{\nu^{n}(X_{1})}{\nu^{n}(X_{2})}\leq\hat{e}(X_{1},X_{2}).$
By pre-compactness, there exists a finite set $F$ such that $\overline{X_{1}}\subset\cup_{y\in F}\mathcal{W}^{u}(y,\delta)$. In particular, using (3.4), one has
$\forall n\in\mathbb{N},\quad\nu^{n}(X_{1})\leq\sum_{y\in F}\nu^{n}(\mathcal{W}^{u}(y,\delta))\leq|F|\times C_{2}(\delta).$
On the other hand, since $X_{2}$ is open, there exists $z\in X_{2}$ and $\delta^{\prime}>0$ such that $\mathcal{W}^{u}(z,\delta^{\prime})\subset X_{2}$. Using (3.4) again, one has
(3.11) $\forall n\in\mathbb{N},\quad C_{1}(\delta^{\prime})\leq\nu^{n}(\mathcal{W}^{u}(z,\delta^{\prime}))\leq\nu^{n}(X_{2}).$
This implies that
$\forall n\in\mathbb{N},\quad\frac{\nu^{n}(X_{1})}{\nu^{n}(X_{2})}\leq\frac{|F|\times C_{2}(\delta)}{C_{1}(\delta^{\prime})}.$
We deduce (3.10) by exchanging the roles of $X_{1}$ and $X_{2}$.
Now, let $\psi\in\mathrm{Con}_{+}^{u}(f)$ and $\phi\in\mathrm{Con}^{u}(f)$ be test functions. There is a $r>0$ such that $A_{r}:=\psi^{-1}(r,+\infty)$ is open and pre-compact. Choose $A$ open and relatively compact containing the support of $\phi$, then using (3.10) for $A_{r}$ and $A$, we obtain
$\forall n\in\mathbb{N},\quad\frac{\nu^{n}(\phi)}{\nu^{n}(\psi)}\leq\frac{\|\phi\|_{\infty}\nu^{n}(A)}{r\nu^{n}(A_{r})}\leq\frac{\|\phi\|_{\infty}}{r}\hat{e}(A,A_{r})=:\hat{e}(\phi,\psi).$
In total, we have shown that:
  1. (1)
    for any $\phi\in\mathrm{Con}^{u}(f)$, there exists a constant $c(\phi)>0$ such that for any $\mu\in\mathcal{X}$, one has $\mu(\phi)\leq c(\phi)$;
  2. (2)
    for any $\phi\in\mathrm{Con}_{+}^{u}(f)$, there exists a constant $c^{\prime}(\phi)>0$ such that for any $\mu\in\mathcal{X}$, one has $\mu(\phi)\geq c^{\prime}(\phi)$.
By Tychonoff’s theorem, this implies that $\mathcal{X}\subset\mathbb{R}^{\mathrm{Con}^{u}(f)}$ is compact for the product topology. ∎
We now construct the $u$-system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ as a fix point. Indeed, define
(3.12) $S:\mathcal{X}\to\mathcal{X},\quad S(\mu):=e^{-h^{u}_{\mathrm{top}}(f)}f^{*}\mu.$
Note that $S$ leaves $\mathcal{X}$ invariant and is continuous. Since $\mathcal{X}$ is convex and compact, the Schauder-Tychonoff fix point Theorem implies that $S$ has a fix point $\tilde{\mu}^{u}$:
(3.13) $\exists\tilde{\mu}^{u}\in\mathcal{X}\quad f^{*}\tilde{\mu}^{u}=e^{h^{u}_{\mathrm{top}}(f)}\tilde{\mu}^{u}$
Since $\tilde{\mu}^{u}\in\mathcal{X}$, (3.4) implies that
(3.14) $\forall x\in\mathbb{T}^{3},\quad C_{1}(\delta)\leq\tilde{\mu}^{u}_{x}(\mathcal{W}^{u}(x,\delta))\leq C_{2}(\delta),$
which shows that $(\tilde{\mu}^{u}_{x})_{x\in\mathbb{T}^{3}}$ has full support in each leaf and is finite on compact sets. Note that it is not clear at this point that $(\tilde{\mu}^{u}_{x})_{x\in\mathbb{T}^{3}}$ is measurable.
A definition of the system of measures using dimension theory. We now define the Margulis system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ using the system of measures $(\tilde{\mu}^{u}_{x})_{x\in\mathbb{T}^{3}}$ constructed above and using ideas from dimension theory, following Climenhaga, Pesin et Zelerowicz [undefk]. This more explicit construction will allow us to show that the system is measurable. We recall the definition of a $u$-Bowen ball:
(3.15) $\forall x\in\mathbb{T}^{3},\ \forall n\geq 0,\ \forall\delta>0,\quad\mathcal{W}^{u}_{n}(x,\delta):=\{y\in\mathbb{T}^{3}\mid\max_{k=0}^{n}d^{u}(f^{k}x,f^{k}y)<\delta\}.$
Since we work with an adapted metric, notice that $\mathcal{W}^{u}_{n}(x,\delta)=f^{-n}\mathcal{W}^{u}(f^{n}(y),\delta)$. For a Borel set $Z\subset\mathcal{W}^{u}(x)$, the set of open covers by $u$-Bowen balls is
(3.16) $\forall N\geq 0,\quad\mathcal{C}_{N}(Z):=\{\mathcal{C}\subset\mathbb{T}^{3}\times\llbracket N,+\infty\llbracket,\ Z\subset\cup_{(y,n)\in\mathcal{C}}\mathcal{W}^{u}_{n}(y,\delta)\}.$
The system of measures is defined as follows:
(3.17) $\mu^{u,N}_{x}(Z):=\inf_{\mathcal{C}\in\mathcal{C}_{N}(Z)}\sum_{(y,n)\in\mathcal{C}}\lambda_{u}^{-n},\quad\mu^{u}_{x}(Z):=\lim_{N\to+\infty}\mu^{u,N}_{x}(Z).$
For any $x\in\mathbb{T}^{3}$, $\mu^{u}_{x}$ defines a Borelian measure on $\mathcal{W}^{u}(x)$. It is clear from (3.17) that $\mu^{u}_{x}(Z)\in[0,+\infty]$ for any Borel subset $Z\subset\mathcal{W}^{u}(x)$. We show that $\mu^{u}_{x}$ is an metric outer measure. Recall that a measure $m$ on a metric space $(X,d)$ is a metric outer measure if for any $A,B\subset X$ such that $d(A,B):=\inf\{d(a,b)\mid a\in A,b\in B\}>0$, one has $m(A\cup B)=m(A)+m(B)$. Every metric outer measure has the property that all Borel sets are measurable, see for instance [undefah, Proposition 12.41].
Suppose that $Z_{1},Z_{2}\subset\mathcal{W}^{u}(x)$ are such that $d^{u}(Z_{1},Z_{2})>0$. Since the $u$-Bowen balls shrink exponentially fast when $N\to+\infty$, for $N$ large enough, $\mathcal{C}_{N}(Z_{1}\cup Z_{2})$ splits as a part that covers $Z_{1}$ and a part that covers $Z_{2}$. In particular, one has
$\forall N\gg 1,\ \mu^{u,N}_{x}(Z_{1}\cup Z_{2})=\mu^{u,N}_{x}(Z_{1})+\mu^{u,N}_{x}(Z_{2})\ \Rightarrow\ \mu^{u}_{x}(Z_{1}\cup Z_{2})=\mu^{u}_{x}(Z_{1})+\mu^{u}_{x}(Z_{2}).$
This shows that $\mu^{u}_{x}$ is a Borel measure on $\mathcal{W}^{u}(x)$.
The measure $\mu^{u}_{x}$ is finite on compact sets. Let $K\subset\mathcal{W}^{u}(x)$ be a compact subset. Let $N\geq 0$ and consider $\mathcal{C}\in\mathcal{C}_{N}(K)$ such that no proper subcover of $\mathcal{C}$ is in $\mathcal{C}_{N}(K)$. One has
$K\subset\bigcup_{(y,n)\in\mathcal{C}}\mathcal{W}^{u}_{n}(y,\delta)=\bigcup_{(y,n)\in{\mathcal{C}}}f^{-n}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big).$
Since $\mathcal{W}^{u}(x)$ is one dimensional and since no proper subcover of $\mathcal{C}$ is in $\mathcal{C}_{N}(K)$, for any $z\in K$, there are at most two elements $(y,n)\in{\mathcal{C}}$ such that $z\in\mathcal{W}^{u}_{n}(y,\delta)$. Hence,
$\sum_{(y,n)\in\mathcal{C}}\tilde{\mu}^{u}_{x}\big(f^{-n}\big(W^{u}(f^{n}y,\delta)\big)\big)\leq 2\tilde{\mu}^{u}_{x}(K).$
Using (3.13) and (3.14), we obtain
$\sum_{(y,n)\in\mathcal{C}}\lambda_{u}^{-n}\leq 2C_{1}(\delta)^{-1}\tilde{\mu}^{u}_{x}(K).$
This shows that $\mu^{u}_{x}(K)\leq 2C_{1}(\delta)^{-1}\tilde{\mu}^{u}_{x}(K)$ is finite.
The measure $\mu^{u}_{x}$ has full support in $\mathcal{W}^{u}(x,\delta)$. Let $x\in\mathbb{T}^{3}$ and $\varepsilon>0$. We show that $\mu^{u}_{x}(\mathcal{W}^{u}(x,\varepsilon))>0$. Let $\mathcal{C}\in\mathcal{C}_{N}(\mathcal{W}^{u}(x,\varepsilon))$. Since $\mathcal{W}^{u}(x,\varepsilon)$ is relatively compact, we can suppose without loss of generality that $\mathcal{C}$ is finite. We have
$\mathcal{W}^{u}(x,\varepsilon)\subset\bigcup_{(y,n)\in\mathcal{C}}\mathcal{W}^{u}_{n}(y,\delta)=\bigcup_{y\in\mathcal{C}}f^{-n}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big).$
This implies that
$\tilde{\mu}^{u}_{x}(\mathcal{W}^{u}(x,\varepsilon))\leq\sum_{(y,n)\in\mathcal{C}}\tilde{\mu}^{u}_{x}\big(f^{-n}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big)\big).$
Using (3.13) and (3.14), we obtain
$C_{1}(\varepsilon)C_{2}(\delta)^{-1}\leq\sum_{(y,n)\in\mathcal{C}}\lambda_{u}^{-n}.$
Passing to the infimum gives $\mu^{u}_{x}(\mathcal{W}^{u}(x,\varepsilon))\geq C_{1}(\varepsilon)C_{2}(\delta)^{-1}>0$.
The Margulis system has the scaling property (3.8). Let $Z\subset\mathcal{W}^{u}(x)$ be a relatively compact Borel set. We need to show that $\mu^{u}_{f(x)}(f(Z))=\lambda_{u}\mu^{u}_{x}(Z)$. Let $N\geq 0$ be an integer and $\mathcal{C}\in C_{N}(Z)$. Then if we define $\mathcal{C}^{\prime}:=\{(f(x),n-1)\mid(x,n)\in\mathcal{C}_{N}(Z)\}$, we see that $\mathcal{C}^{\prime}\in\mathcal{C}_{N-1}(f(Z))$. Moreover, any $\mathcal{C}^{\prime}\in\mathcal{C}_{N-1}(f(Z))$ can be obtained this way and
$\sum_{(y,m)\in\mathcal{C}^{\prime}}\lambda_{u}^{-m}=\sum_{(z,n)\in\mathcal{C}}\lambda_{u}^{-(n-1)}=\lambda_{u}\sum_{(z,n)\in\mathcal{C}}\lambda_{u}^{-n}.$
Taking the infimum yields $\mu^{u}_{f(x)}(f(Z))=\lambda_{u}\mu^{u}_{x}(Z)$.
The Margulis system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ is measurable. Since the limit of a sequence of measurable functions is measurable, (3.17) shows that it suffices to prove that $(\mu^{u,N}_{x})_{x\in\mathbb{T}^{3}}$ is measurable. Let $M\in\mathbb{N}$ with $M>N$. Define
$\mathcal{C}_{N,M}(Z):=\{\mathcal{C}\subset\mathbb{T}^{3}\times\llbracket N,M\rrbracket,\ Z\subset\cup_{(y,n)\in\mathcal{C}}\mathcal{W}^{u}_{n}(y,\delta)\}$
as well as
(3.18) $\mu^{u,N,M}_{x}(Z):=\inf_{\mathcal{C}\in\mathcal{C}_{N,M}(Z)}\sum_{(x,n)\in\mathcal{C}}\lambda_{u}^{-n}.$
Since $\mu^{u,N}_{x}:=\lim_{M\to+\infty}\mu^{u,N,M}_{x}$, we only need to show that $(\mu^{u,N,M}_{x})_{x\in\mathbb{T}^{3}}$ is measurable. To that end, first consider a compact set $K\subset\mathbb{T}^{3}$. We prove that $x\mapsto\mu^{u,N,M}_{x}(K_{x})$, where $K_{x}:=K\cap\mathcal{W}^{u}_{\mathrm{loc}}(x)$, is upper semi-continuous, thus measurable. By compactness of $K$, $x\mapsto K_{x}$ is upper semicontinuous. For our purpose, this means that for any $x\in\mathbb{T}^{3}$ and any $\varepsilon>0$, there is a $\eta>0$ such that if $d(x,y)<\eta$, then
$K_{y}\subset\mathrm{Hol}^{cs}_{x,y}(U_{\varepsilon}(K_{x}))$
where $U_{\varepsilon}(K_{x}):=\cup_{z\in K_{x}}\mathcal{W}^{u}(z,\varepsilon)$ is the $\varepsilon$-neighborhood of $K_{x}$. Let $\mathcal{C}\in\mathcal{C}_{N,M}(K_{x})$ with no proper subcover in $\mathcal{C}_{N,M}(K_{x})$. This means that
$K_{x}\subset\bigcup_{(y,n)\in\mathcal{C}}f^{-n}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big).$
As before, this implies
$\sum_{(y,n)\in\mathcal{C}}\tilde{\mu}^{u}_{x}\big(f^{-M}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big)\big)\leq\sum_{(y,n)\in\mathcal{C}}\tilde{\mu}^{u}_{x}\big(f^{-n}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big)\big)\leq 2\tilde{\mu}^{u}_{x}(K).$
Using (3.13) and (3.14), we obtain $|\mathcal{C}|\leq 2C_{1}(\delta)^{-1}\tilde{\mu}^{u}_{x}(K_{x})\lambda_{u}^{M}.$ In particular, the infimum in (3.18) can be taken over partitions of uniformly bounded cardinality. As a consequence, the function in (3.18) takes its values in a finite set and thus attains its minimum. We now suppose that $\mathcal{C}\in\mathcal{C}_{N,M}(K_{x})$ is such that
$\mu^{u,N,M}_{x}(K_{x})=\sum_{(y,n)\in\mathcal{C}}\lambda_{u}^{-n}.$
Since $K_{x}$ is closed, there exists $\varepsilon>0$ and $\delta^{\prime}<\delta$ such that
$U_{\varepsilon}(K_{x})\subset\bigcup_{(y,n)\in\mathcal{C}}f^{-n}\big(W^{u}(f^{n}y,\delta^{\prime})\big).$
Let $y\in\mathbb{T}^{3}$ be $\eta$-close to $x$ where $\eta$ is defined by the upper semi-continuity of $K$. Up to taking $\eta$ even smaller and since all the foliations are continuous, we have
$\mathrm{Hol}^{cs}_{f^{n}x,f^{n}y}\big(W^{u}(f^{n}z,\delta^{\prime})\big)\subset\mathcal{W}^{u}(\mathrm{Hol}^{cs}_{f^{n}x,f^{n}y}\circ f^{n}z,\delta)=\mathcal{W}^{u}(f^{n}\circ\mathrm{Hol}^{cs}_{x,y}(z),\delta).$
In other words, we have shown that $\mathcal{C}^{\prime}:=\{(\mathrm{Hol}^{cs}_{x,y}(z),n)\mid(z,n)\in\mathcal{C}\}$ is an element of $\mathcal{C}_{N,M}(K_{y})$. Moreover, one has
$\mu^{u,N,M}_{x}(K_{x})=\sum_{(z,n)\in\mathcal{C}}\lambda_{u}^{-n}=\sum_{(w,n)\in\mathcal{C}^{\prime}}\lambda_{u}^{-n}.$
Passing to the infimum, we thus have $\mu^{u,N,M}_{y}(K_{y})\leq\mu_{x}^{u,N,M}(K_{x})$ which finishes the proof for compact sets $K$.
Now, let $\mathcal{K}$ be the collection of all subsets $Z\subset\mathbb{T}^{3}$ such that $x\mapsto\mu^{u,N,M}_{x}(Z_{x})$ is measurable. Note that this is a Dynkin system, i.e.,
  • If $Z_{1},Z_{2}\in\mathcal{K}$ and $Z_{1}\subset Z_{2}$, then $Z_{2}\setminus Z_{1}\in\mathcal{K}$.
  • If $Z_{1}\subset Z_{2}\subset\ldots\subset Z_{n}\subset\ldots\in\mathcal{K}$, then $\cup_{n\in\mathbb{N}}Z_{n}\in\mathcal{K}$.
  • $\mathbb{T}^{3}\in\mathcal{K}$.
Since $\mathcal{K}$ contains the closed subsets, Dynkin $\pi$-$\lambda$ theorem implies that it must contain the $\sigma$-algebra generated by the closed subsets. In other words, $\mathcal{K}$ contains all Borel sets which concludes the proof. ∎
Having constructed a measurable system of $u$-measures, we now characterize $u$-MMEs by their conditionals along the strong unstable foliation. We recall the following result due to Buzzi, Fisher and Tazhibi [undefh], see also [undefam, Proposition 5.3].
Proposition 3.6.
Let $f\in\mathrm{Diff}^{\infty}(M)$ be a partially hyperbolic diffeomorphism on a closed manifold $M$ with a measurable Margulis system of measures $(\nu^{u}_{x})_{x\in M}$ such that
  • for any $x\in M$, $\nu^{u}_{x}$ is fully supported in $\mathcal{W}^{u}(x)$,
  • there is $D_{u}>0$ such that for any $x\in M$, $f_{*}\nu^{u}_{x}=D_{u}\nu^{u}_{f(x)}.$
Then for any invariant measure $\nu$, one has $h^{u}_{\nu}(f)\leq\ln(D_{u}),$ with equality if and only if the disintegration of $\nu$ along $\mathcal{W}^{u}$ is given by $\nu^{u}_{x}$ ((up to a constant rescaling)), $\nu$-a.e.
Using Corollary 3.2 and Proposition 3.4, we deduce Corollary 1.2.
4. The first Ruelle resonance on $2$-forms.
In this section, we study the first resonance of the Anosov diffeomorphism $f$ when acting on the bundle of differential $2$-forms. We will show the following result. Recall that the space of resonant states $\mathrm{Res}_{k}(f)$ is defined in (1.6) and that $\lambda\in\mathrm{Res}_{k}(f)\cap\mathbb{R}_{+}$ is said to be the first resonance for the action on $k$-forms if $\mathrm{Res}_{k}(f)\subset\{z\in\mathbb{C},|z|\leq\lambda\}.$
Proposition 4.1 (First Ruelle resonance).
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. Then one has
(4.1) $\sup\{|z|,z\in\mathrm{Res}_{2}(f)\}=\lambda_{u}=e^{h_{\mathrm{top}}^{u}(f)}=e^{P(J^{c}_{f})}.$
Moreover, $\lambda_{u}$ is the first Pollicott-Ruelle resonance, it has no Jordan block and has multiplicity equal to one. Finally, the trace of the spectral projector of $f$ at $z=\lambda_{u}$ is equal to $\mu_{J^{c}_{f}}$.
We note that as a direct consequence of (4.1) and (2.13), we obtain.
Corollary 4.2.
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. The equilibrium state $\mu_{J^{c}_{f}}$ is a $u$-MME.
4.1. Ruelle resonances
In this subsection, we recall some facts about the functional approach for Anosov diffeomorphisms. For a detailed introduction to the subject, we refer to [undefa]. For $0\leq k\leq 3$, define the vector bundle of smooth $k$-forms $\Omega_{k}=C^{\infty}(\mathbb{T}^{3};\Lambda^{k}T^{*}(\mathbb{T}^{3})).$ The diffeomorphism $f$ acts on $\Omega_{k}$ by pushforward:
(4.2) $f_{*,k}:\Omega_{k}\to\Omega_{k},\quad f_{*,k}\omega:=f_{*}\omega.$
For an Anosov diffeomorphism, one can associate to $f_{*,k}$ a discrete spectrum $\mathrm{Res}_{k}(f)$, the Pollicott-Ruelle resonances, by making it act on specially designed anisotropic spaces, see for instance [undefd, undefb, undefq, undefo].
For any $0\leq k\leq 3$, the resolvent $R_{k}(\lambda):=(f_{*,k}-\lambda)^{-1}:L^{2}(\mathbb{T}^{3};\Omega_{k})\to L^{2}(\mathbb{T}^{3};\Omega_{k}),$ defined for $\lambda\in\mathbb{C}$ such that $|\lambda|\gg 1$ admits a meromorphic extension $R_{k}(\lambda)$ to $\mathbb{C}\setminus\{0\}$
(4.3) $R_{k}(\lambda):\Omega_{k}\to\mathcal{D}^{\prime}_{k}(\mathbb{T}^{3}),$
where $D^{\prime}_{k}(\mathbb{T}^{3})$ denotes the space of distributional currents of degree $k$, i.e., the dual space of $\Omega_{3-k}$, see [undefab, Chapter 10.1]. We will write $(\cdot,\cdot)_{\Omega_{k} \times\mathcal{D}^{\prime}_{3-k}}$ for the pairing between $\Omega_{k}$ and $\mathcal{D}^{\prime}_{3-k}$. More precisely, by the work of Faure, Roy and Sjöstrand [undefo], there exists a family of Hilbert spaces $(\mathcal{H}^{s}(\mathbb{T}^{3};\Omega_{k}))_{s>0}$ such that:
  • the space $\Omega_{k}$ is densely included in $\mathcal{H}^{s}(\mathbb{T}^{3};\Omega_{k})$ (see [undefab, Lemma 9.1.13]) and one has $H^{s}\subset\mathcal{H}^{s}\subset H^{-s}$, where $H^{s}$ is the usual $L^{2}$-Sobolev space of order $s$ (see [undefab, Lemma 9.1.14]);
  • there exists $c>0$ such that for any $s>0$, the resolvent $R_{k}(\lambda)=\mathcal{H}^{s}\to\mathcal{H}^{s}$ is well defined, bounded and holomorphic for $\{|\lambda|\gg 1\}$, and has a meromorphic extension to $\{|\lambda|>e^{-cs}\}$ independent of any choice made in the construction.
Applying the result to $f^{-1}$, we obtain a family of Hilbert spaces $(\mathcal{H}^{-s}(\mathbb{T}^{3},\Omega_{k}))_{s>0}$, dual to $(\mathcal{H}^{s}(\mathbb{T}^{3},\Omega_{k}))_{s>0}$ on which the resolvent of the pullback of $f$ has a meromorphic extension, see [undefo]. We will write $(\cdot,\cdot)_{\mathcal{H}^{s} \times\mathcal{H}^{-s}}$ for the pairing between $\mathcal{H}^{s}$ and $\mathcal{H}^{-s}$.
The poles of the extension are intrinsic and are called the Pollicott-Ruelle resonances $\mathrm{Res}_{k}(f)$ of $f_{*,k}$. For any $\lambda_{0}\in\mathrm{Res}_{k}(f)$, the spectral projector at $\lambda_{0}$ is given by
$\Pi_{k}^{\lambda_{0}}=-\frac{1}{2i\pi}\int_{\gamma}R_{k}(z)dz,$
where $\gamma$ is a small loop around $\lambda_{0}$ and has finite rank. The (algebraic) multiplicity of $\lambda_{0}$ is given by the rank of the spectral projector:
(4.4) $\forall\lambda_{0}\in\mathrm{Res}_{k}(f),\quad m_{k}(\lambda_{0}):=\mathrm{rank}(\Pi_{k}^{\lambda_{0}})\in\mathbb{N}.$
Near a resonance $\lambda_{0}\in\mathrm{Res}_{k}(f)$, the resolvent admits a Laurent expansion of the form
$R_{k}(\lambda)=R_{k}^{H}(\lambda)-\sum_{j=1}^{N(\lambda_{0})}\frac{(f_{*,k}-\lambda_{0})^{j-1}\Pi_{\lambda_{0}}^{+}}{(\lambda-\lambda_{0})^{j}},$
where $R_{k}^{H}(\lambda)$ is the holomorphic part near $\lambda_{0}$. The generalized resonant states are the elements in the range of the spectral projector.
$\mathrm{Res}_{k,\lambda_{0},\infty}(f):=\Pi_{\lambda_{0}}^{+}(\mathcal{H}^{s})=\Pi_{\lambda_{0}}^{+}(\Omega_{k})=\{u\in\mathcal{H}^{s}\mid(f_{*,k}-\lambda_{0})^{N(\lambda_{0})}u=0\}.$
Since the meromorphic extension is independent of the particular choice of anisotropic space $\mathcal{H}^{s}$, the space of generalized resonant states does not depend on $\mathcal{H}^{s}$. One way to see this is to use the following equivalent characterization of generalized resonant states. Define a dual decomposition of the Anosov decomposition:
(4.5) $T^{*}(\mathbb{T}^{3})=E_{s}^{*}\oplus E_{cu}^{*}=E_{cs}^{*}\oplus E_{u}^{*},\quad E_{\bullet}^{*}(E_{\bullet})=0,\ \bullet=s,cu,cs,u.$
Then one has, see [undefn, Lemma 5.12],
$\mathrm{Res}_{k,\lambda_{0},\infty}(f)=\{u\in\mathcal{D}^{\prime}_{k}(\mathbb{T}^{3})\mid(f_{*,k}-\lambda_{0})^{N(\lambda_{0})}u=0,\ \mathrm{WF}(u)\subset E_{cu}^{*}\},$
where $\mathrm{WF}(u)$ denotes the wavefront set of a current $u$, see [undefv, Chapter VIII]. The resonant states are defined as
(4.6) $\mathrm{Res}_{k,\lambda_{0}}(f):=\{u\in\mathcal{D}^{\prime}_{k}(\mathbb{T}^{3})\mid(f_{*,k}-\lambda_{0})u=0,\ \mathrm{WF}(u)\subset E_{cu}^{*}\}.$
We will say that $f$ has no Jordan block at $\lambda_{0}$ if $N(\lambda_{0})=1$, that is, if $\mathrm{Res}_{k,\lambda_{0}}(f)=\mathrm{Res}_{k,\lambda_{0},\infty}(f)$. Note that applying the previous results to $f^{-1}$ yields notions of (generalized) co-resonant states, see (1.7).
4.2. Zeta function
We recall the definition of the Ruelle zeta function. In our setting, this function is also known as the Artin-Mazur zeta function. For $n\geq 1$, let
(4.7) $\mathcal{F}_{n}(f):=\{x\in\mathbb{T}^{3}\mid f^{n}(x)=x\}.$
The Ruelle zeta function is given by:
(4.8) $\forall z\in\mathbb{C},\ |z|\ll 1,\quad\zeta_{f}(z):=\exp\left(\sum_{m=1}^{+\infty}\frac{z^{m}}{m}\mathrm{Card}(\mathcal{F}_{m}(f))\right).$
The zeta function $\zeta_{f}$ has a meromorphic extension (actually rational extension in our case) to the complex plane. This is a special case of the works of Liverani-Tsujii [undefac] and Baladi-Tsujii [undefc]. In fact, for an Anosov diffeomorphism on a torus, one can compute explicitly $\zeta_{f}$ in terms of its action on the homology $A$.
Lemma 4.3.
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ and let $A\in\mathrm{GL}_{3}(\mathbb{Z})$ be its action on homology. Then
(4.9) $\forall z\in\mathbb{C},\quad\zeta_{f}(z)=\zeta_{A}(z)=\frac{(1-z\lambda_{s}\lambda_{c})(1-z\lambda_{s}\lambda_{u})(1-z\lambda_{c}\lambda_{u})}{(1-z\lambda_{s})(1-z\lambda_{c})(1-z\lambda_{u})}.$
Proof.
Since $f$ preserves orientation, the Lefschetz fixed point Theorem gives
$\mathrm{Card}\big(\mathcal{F}_{m}(f)\big)=\sum_{k=0}^{3}(-1)^{k}\operatorname{tr}(\tilde{f}_{k,*}^{m}),$
where $\tilde{f}_{k,*}:H_{k}(\mathbb{T}^{3},\mathbb{R})\cong\mathbb{Z}^{\binom{3}{k}}\to H_{k}(\mathbb{T}^{3},\mathbb{R})$ denotes the action of $f$ on the $k$-th homology group of $\mathbb{T}^{3}$. Since $\tilde{f}_{k,*}=\tilde{A}_{k,*}$, an easy computation using (4.8) yields
$\zeta_{f}(z)=\zeta_{A}(z)=\prod_{k=0}^{3}\mathrm{det}(\mathrm{Id}-z\tilde{A}_{k,*})^{(-1)^{k}}=\frac{(1-z)(1-z\lambda_{s}\lambda_{c})(1-z\lambda_{s}\lambda_{u})(1-z\lambda_{c}\lambda_{u})}{(1-z)(1-z\lambda_{s})(1-z\lambda_{c})(1-z\lambda_{u})},$
which simplifies into (4.9). ∎
4.3. Dynamical determinants
In this subsection, we recall the link between the spectral theory of the diffeomorphism $f$ on the anisotropic spaces and the periodic orbits of $f$. We define the dynamical determinants. For $k\in\llbracket 0,3\rrbracket$, let
(4.10) $\forall z\in\mathbb{C},|z|\ll 1,\quad\mathcal{D}_{f,k}(z):=\exp\left(-\sum_{m=1}^{+\infty}\frac{z^{m}}{m}\sum_{x\in\mathcal{F}_{m}(f)}\frac{\operatorname{tr}(\Lambda^{k}df^{-m}(x))}{|\mathrm{det}(\mathrm{Id}-df^{-m}(x))|}\right).$
We will need the following result, see for instance [undefa, Theorem 6.2]. For any $k\in\llbracket 0,3\rrbracket$, the dynamical determinant $\mathcal{D}_{f,k}$ admits a holomorphic extension to $\mathbb{C}$. Moreover,
(4.11) $\forall z\in\mathbb{C}^{*},\quad\mathcal{D}_{f,k}(z)=0\ \iff\ z^{-1}\in\mathrm{Res}_{k}(f),$
that is, $z$ is a zero of $\mathcal{D}_{f,k}$ if and only if $z^{-1}$ is a Pollicott-Ruelle resonance of $f$ for the action on $k$-forms. Moreover, the multiplicity of the zero coincides with the algebraic multiplicity of $z^{-1}$. Using (4.11), we now show the following lemma.
Lemma 4.4.
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, then one has
$\sup\{|z|,z\in\mathrm{Res}_{2}(f)\}=e^{P(J^{c}_{f})},\quad e^{P(J^{c}_{f})}\in\mathrm{Res}_{2}(f).$
We will say that $e^{P(J^{c}_{f})}$ is the first resonance of $f$ for the action on $2$-forms. Moreover, $e^{P(J^{c}_{f})}$ has no Jordan block, has multiplicity $1$ as a Pollicott-Ruelle resonance and it is the only resonance on the critical circle $\{z\in\mathbb{C},|z|=e^{P(J^{c}_{f})}\}$.
Proof.
From (4.11), it suffices to locate the first zero of $\mathcal{D}_{f,2}$ and compute its multiplicity. Let $x\in\mathbb{T}^{3}$ and let $n\geq 1$. Since $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, the linear map $df^{-m}(x)$ has three eigenvalues $\rho^{m}_{\bullet}(x):=\mathrm{det}(df^{-m}(x)|_{E_{\bullet}(x)})$ for $\bullet=s,c,u.$ Moreover, there exists constants $C,\nu>0$ such that for any $x\in\mathbb{T}^{3}$ and $m\geq 1$,
(4.12) $\rho^{m}_{s}(x)\geq Ce^{\nu m}\geq 1\geq Ce^{-\nu m}\geq\rho^{m}_{c}(x)\geq Ce^{-\nu m}\rho^{m}_{c}(x)\geq\rho^{m}_{u}(x).$
In particular, we deduce that, when $m\to+\infty$,
(4.13) $\begin{split}&|\mathrm{det}(\mathrm{Id}-df^{-m}(x))|=\rho^{m}_{s}(x)\big(1+O(e^{-\nu m})\big),\\ &\operatorname{tr}(\Lambda^{2}df^{-m}(x))=\rho^{m}_{s}(x)\rho^{m}_{c}(x)\big(1+O(e^{-\nu m})\big),\end{split}$
uniformly in $x\in\mathbb{T}^{3}$. Plugging (4.13) into (4.10) gives
$\sum_{x\in\mathcal{F}_{m}(f)}\frac{\operatorname{tr}(\Lambda^{2}df^{-m}(x))}{|\mathrm{det}(\mathrm{Id}-df^{-m}(x))|}=\sum_{x\in\mathcal{F}_{m}(f)}\rho_{c}^{m}(x)\big(1+O(e^{-\nu m})\big)\sim_{m\to+\infty}\sum_{x\in\mathcal{F}_{m}(f)}\rho_{c}^{m}(x).$
Now, using the chain rule and (2.9), we obtain
$\forall m\geq 1,\ \forall x\in\mathcal{F}_{m}(f),\quad\rho^{m}_{c}(x)=\mathrm{det}(df^{m}(f^{-m}x)|_{E_{c}(x)})^{-1}=\exp\left(\sum_{k=0}^{m-1}J_{f}^{c}(f^{k}x)\right).$
Since $\rho^{m}_{c}(x)>0$, the zeros of $\mathcal{D}_{f,2}$ on the critical circle are the same as the zeros of
$\zeta_{f,J^{c}_{f}}(z):=\exp\left(-\sum_{m=1}^{+\infty}\frac{z^{m}}{m}\sum_{x\in\mathcal{F}_{m}(f)}\exp\left(\sum_{k=0}^{m-1}J_{f}^{c}(f^{k}x)\right)\right).$
We recognize the Ruelle zeta function with potential $J^{c}_{f}$. Since $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is conjugate to a linear automorphism of $\mathbb{T}^{3}$, it is topologically mixing. By [undefai, Theorem 1], there is a $R=R(J^{c}_{f})>e^{-P(J^{c}_{f})}$ such that $\zeta_{f,J^{c}_{f}}$ has no zero in $\overline{B}(0,R)$ except for a simple zero at $e^{-P(J^{c}_{f})}$. This shows that $e^{P(J^{c}_{f})}$ is a resonance of algebraic multiplicity equal to $1$ and that $e^{P(J^{c}_{f})}$ is the only resonance on the critical circle. ∎
The dynamical determinants can be linked to the zeta function,
(4.14) $\zeta_{f}(z)=\prod_{k=0}^{3}\mathcal{D}_{f,k}(z)^{(-1)^{k+1}}=\frac{\mathcal{D}_{f,1}(z)\mathcal{D}_{f,3}(z)}{\mathcal{D}_{f,0}(z)\mathcal{D}_{f,2}(z)},$
see [undefa, Equation $(6.38)$]. Using (4.9), we also show the following.
Lemma 4.5.
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. Then
$\sup\{|z|,z\in\mathrm{Res}_{2}(f)\}=\lambda_{u}.$
Proof.
Combining (4.14) and (4.9), we get
(4.15) $\frac{\mathcal{D}_{f,1}(z)\mathcal{D}_{f,3}(z)}{\mathcal{D}_{f,0}(z)\mathcal{D}_{f,2}(z)}=\frac{(1-z\lambda_{s}\lambda_{c})(1-z\lambda_{s}\lambda_{u})(1-z\lambda_{c}\lambda_{u})}{(1-z\lambda_{s})(1-z\lambda_{c})(1-z\lambda_{u})}.$
Since the dynamical determinants are holomorphic, the pole $\lambda_{u}^{-1}$ can only come from either a zero of $\mathcal{D}_{f,0}$ or a zero of $\mathcal{D}_{f,2}$. In other words, we have either $\lambda_{u}\in\mathrm{Res}_{0}(f)$ or $\lambda_{u}\in\mathrm{Res}_{2}(f)$. Now, for any $g\in C^{\infty}(\mathbb{T}^{3})$, one has $\|f_{*}g\|_{\infty}\leq\|g\|_{\infty}$. Hence, for any $z\in\mathbb{C}$ such that $|z|>1$, the resolvent $(f_{*}-z)^{-1}:L^{\infty}(\mathbb{T}^{3})\to L^{\infty}(\mathbb{T}^{3})$ is defined by
$(f_{*}-z)^{-1}g=z^{-1}\sum_{m=0}^{+\infty}\frac{f^{n}_{*}g}{z^{n}}.$
In particular, this shows that $\mathrm{Res}_{0}(f)\subset\{z \in\mathbb{C},\ |z|\leq 1\}$ and thus $\lambda_{u}\notin\mathrm{Res}_{0}(f)$. We deduce that $\lambda_{u}\in\mathrm{Res}_{2}(f)$ is a Pollicott-Ruelle resonance. Using Lemma 4.4, this shows that $\lambda_{u}\leq e^{P(J^{c}_{f})}$. By Corollary 3.2, this means that $h^{u}_{\mathrm{top}}(f)\leq P(J^{c}_{f})$. Using (2.11) we deduce $h^{u}_{\mathrm{top}}(f)=P(J^{c}_{f})$, which concludes the proof of the lemma and of (4.1). ∎
We show that the spectral projector at the first resonance is equal to $\mu_{J^{c}_{f}}$.
Lemma 4.6.
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. The trace of the spectral projector at the first resonance $e^{P(J^{c}_{f})}$ is given by the product of a resonant state $\theta$ and a co-resonant state $\nu$ and is equal to the equilibrium state $\mu_{J^{c}_{f}}$.
Proof.
Let $g\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}_{+})$ be a smooth and positive function. The Guillemin trace formula (see for instance [undefc, Proposition 6.3]) gives for any $m\geq 0$,
$\operatorname{tr}^{\flat}\big(gf_{*,2}^{m}\big)=\sum_{x\in\mathcal{F}_{m}(f)}\frac{g(x)\operatorname{tr}(\Lambda^{2}df^{-m}(x))}{|\mathrm{det}(\mathrm{Id}-df^{-m}(x))|},$
where $\operatorname{tr}^{\flat}$ denotes the flat trace and $f_{*,2}$ is defined in (4.2). For any $z\in\mathbb{C}$, such that $|z|\ll 1$, define $Z_{g,f}(z):=\operatorname{tr}^{\flat}\big(g(z-f_{*,2})^{-1}\big)$. For $|z|\ll 1$,
(4.16) $Z_{g,f}(z)=\sum_{m=0}^{+\infty}z^{-m-1}\operatorname{tr}^{\flat}\big(gf_{*,2}^{m}\big)=\sum_{m=0}^{+\infty}z^{-m-1}\sum_{x\in\mathcal{F}_{m}(f)}\frac{g(x)\operatorname{tr}(\Lambda^{2}df^{-m}(x))}{|\mathrm{det}(\mathrm{Id}-df^{-m}(x))|}.$
Note that $Z_{g,f}$ admits a meromorphic extension to $\mathbb{C}\setminus\{0\}$ since the resolvent does. Using Lemma 4.4, we see that $Z_{g,f}$ has a first simple pole at $z=\lambda_{u}$ and no other pole outside of $B(0,\lambda_{u}^{-1}-\varepsilon)$ for some $\varepsilon>0$. We use Lemma 4.4 and write the spectral projector at the first resonance $\lambda_{u}$ as $\Pi_{0}(\lambda_{u})=\nu(\cdot)\theta,$ where $\theta$ (resp. $\nu$) is a resonant (resp. co-resonant state) which are normalized so that $(\nu,\theta)_{\mathcal{H}^{-s}\times\mathcal{H}^{s}}=1$. Recall that near $z=\lambda_{u}^{-1}$,
$\operatorname{tr}^{\flat}\big(g(z-f_{*,2})^{-1}\big)$ $=-\operatorname{tr}^{\flat}\big(g(f_{*,2}-z)^{-1}\big)=\operatorname{tr}^{\flat}\Big(g\big(\tfrac{\Pi_{0}}{z-\lambda_{u}}+h(z)\big)\Big),$
for some holomorphic map $h$. In particular, the residue at $\lambda_{u}$ is given by
$\mathrm{Res}_{\lambda_{u}}(Z_{g,f})$ $=\operatorname{tr}(g\Pi_{0}(\lambda_{u}))=\nu(g\times\theta)=(\nu,g\theta)_{\mathcal{H}^{-s}\times\mathcal{H}^{s}}=:(\nu\wedge\theta)(g).$
Next, we use Cauchy’s formula on (4.16) to obtain, for some $\rho<\lambda_{u}^{-1}$,
$\sum_{x\in\mathcal{F}_{m}(f)}\frac{g(x)\operatorname{tr}(\Lambda^{2}df^{-m}(x))}{|\mathrm{det}(\mathrm{Id}-df^{-m}(x))|}=$ $\mathrm{Res}_{\lambda_{u}}(Z_{g,f}(z)z^{m+1})+\frac{1}{2i\pi}\int_{\rho\mathbb{S}^{1}}Z_{g,f}(z)z^{m-1}dz.$
Since $Z_{g,f}$ has a simple first pole $\lambda_{u}$, one has
$\sum_{x\in\mathcal{F}_{m}(f)}\frac{g(x)\operatorname{tr}(\Lambda^{2}df^{-m}(x))}{|\mathrm{det}(\mathrm{Id}-df^{-m}(x))|}=(\nu\wedge\theta)(g)\lambda_{u}^{m}(1+O(e^{-\varepsilon m})),$
for some $\varepsilon>0$. Using (4.13), this gives
(4.17) $(\nu\wedge\theta)(g)=\lim_{m\to+\infty}\lambda_{u}^{-m}\sum_{x\in\mathcal{F}_{m}(f)}\exp\left(-\sum_{k=0}^{m-1}J_{f}^{c}(f^{k}x)\right)g(x).$
Using Lemmas 4.4 and 4.6, for $g\equiv 1$, we obtain
$\lim_{m\to+\infty}$ $\lambda_{u}^{m}\sum_{x\in\mathcal{F}_{m}(f)}\mathrm{det}(df^{m}(x)|_{E_{c}(x)})=\lim_{m\to+\infty}e^{mP(J^{c}_{f})}\sum_{x\in\mathcal{F}_{m}(f)}\mathrm{det}(df^{m}(x)|_{E_{c}(x)})=1.$
This means that (4.17) is Bowen’s formula for the equilibrium state (see for instance $\mu_{J^{c}_{f}}$ [undeft, Theorem 20.3.7]) and this concludes the proof. ∎
Recall that by Proposition 4.1, we know that the space of (co)-resonant states at $\lambda_{u}$ is one dimensional and thus spanned by a (co)-resonant state $\theta$ and $\nu$.
Lemma 4.7.
For any $s>0$ large enough and for any $\omega\in\mathcal{H}^{s}$, one has
(4.18) $\exists c_{\omega}\in\mathbb{R},\quad\lim_{n\to+\infty}\lambda_{u}^{-n}(f_{*})^{n}\omega=c_{\omega}\theta,$
where the limit exists in the anisotropic space $\mathcal{H}^{s}$. Similarly, for $\eta\in\mathcal{H}^{-s}$, one has
(4.19) $\exists c_{\eta}\in\mathbb{R},\quad\lim_{n\to+\infty}\lambda_{u}^{-n}(f^{*})^{n}\eta=c_{\eta}\nu,$
where the limit exists in the anisotropic space $\mathcal{H}^{-s}$.
Proof.
By Lemma 4.4, there is a $R>0$ such that $\lambda_{u}$ is the only Pollicott-Ruelle resonance in $\mathbb{C}\setminus B(0,R)$ and it is simple with no Jordan block. We fix $s$ such that $e^{-cs}<R$, so that $R_{k}(\lambda)=\mathcal{H}^{s}\to\mathcal{H}^{s}$ has a meromorphic extension to $\{|\lambda|>R\}$. In particular, denoting by $\Pi_{0}:\mathcal{H}^{s}\to\mathcal{H}^{s}$ the spectral projector on $\mathrm{Span}(\theta)$, there exists a bounded operator $K:\mathcal{H}^{s}\to\mathcal{H}^{s}$ satisfying
$f_{*,k}=\lambda_{u}\Pi_{0}+K,\quad K\Pi_{0}=\Pi_{0}K=0,\quad r(K)<\lambda_{u},$
where $r(K)$ is the spectral radius of $K$. For any $\varepsilon>0$, recall that one has
$\exists C>0,\quad\forall n\geq 0,\quad\|K^{n}\|_{\mathcal{H}^{s}\to\mathcal{H}^{s}}\leq C(r(K)+\varepsilon)^{n}.$
If $\varepsilon>0$ is chosen so that $r(K)+\varepsilon<\lambda_{u}$, then, as bounded operators $\mathcal{H}^{s}\to\mathcal{H}^{s}$,
$\lambda_{u}^{-n}f_{*,k}^{n}=\lambda_{u}^{-n}(\lambda_{u}\Pi_{0}+K)^{n}\to_{n\to+\infty}\Pi_{0}.$
Let $\omega\in\mathcal{H}^{s}$, then there is $c_{\omega}\in\mathbb{R}$ such that $\Pi_{0}(\omega)=c_{\omega}\theta$. We obtain (4.18) by applying the previous convergence to $\omega$. The proof of (4.19) is analogous.
Remark 2.
The resonant ((resp. co-resonant)) state $\theta$ ((resp. $\nu)$ is closed. Indeed, since $d$ commutes with the pushforward (resp. pullback) and preserves the spaces of currents with wavefront set in $E_{s}^{*}$ ((resp. $E_{u}^{*})$, one has $d\theta\in\mathrm{Res}_{3,\lambda_{u}}(f)$ ((resp. $d\nu\in\mathrm{Res}^{*}_{2,\lambda_{u}}(f))$. A similar argument as in the proof of Lemma 4.5 shows that $\mathrm{Res}_{3}(f)\subset\{z\in\mathbb{C},\ |z|\leq 1\}$ which then implies that $d\theta=0$. For the co-resonant state $\nu$, we can use (4.15) with (4.11) to conclude that $\mathrm{dim}(\mathrm{Res}^{*}_{1,\infty,\lambda_{u}}(f))-\mathrm{dim}(\mathrm{Res}^{*}_{2,\infty,\lambda_{u}}(f))=1$. Since Lemma 4.4 shows that $\mathrm{dim}(\mathrm{Res}^{*}_{1,\infty,\lambda_{u}}(f))=1$, we deduce that $\mathrm{Res}^{*}_{2,\infty,\lambda_{u}}(f))=\{0\}$ and thus $d\nu=0$. This should be compared with the work of Ruelle and Sullivan on the system of Margulis measures for an Anosov diffeomorphism, see [undefaj, Corollary p.326].
4.4. Measure of maximal $u$-entropy which is not a measure of maximal entropy.
In this subsection, we show Corollary 1.1.
Proof.
By [undeft, Theorem 20.3.9], $\mu_{J^{c}_{f}}$ is a measure of maximal entropy if and only if $J^{c}_{f}$ is cohomologous to a constant. By the Livsic theorem [undeft, Theorem 19.2.1], this holds if and only if for any periodic point $p$ of period $n(p)$, one has
$\frac{1}{n(p)}\sum_{k=0}^{n(p)-1}J^{c}_{f}(f^{k}p)=\lambda_{c}.$
The fact that the constant has to be equal to $\lambda_{c}$ follows from [undefp, Corollary 2.6]. This holds if and only if $E_{u}$ and $E_{s}$ are jointly integrable by [undefp, Theorem 5.1]. ∎
5. System of Margulis measures as (co)-resonant states.
In this section, we show that the resonant (resp. co-resonant) state at the first resonance $\lambda_{u}$ is given by the system of center stable (resp. strong unstable) measures. The main technical result of this section is that the system of $u$-measure is invariant with respect to the center stable holonomies.
Proposition 5.1.
Let $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. Let $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ be the system of $u$-measures constructed in Theorem 3.4. Let $R$ be a rectangle. For any $x,z\in R$, one has
(5.1) $(\mathrm{Hol}^{cs}_{x,z})_{*}\mu^{u}_{z}=\mu^{u}_{x}.$
5.1. The resonant state defines a system of center stable measures.
We note that by Corollary 1.2 and Corollary 4.2, the system of $u$-measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ coincides (up to a constant rescaling) with the conditionals $(\mu^{u}_{J^{c}_{f},x})_{x\in\mathbb{T}^{3}}$ of $\mu_{J^{c}_{f}}$ along $\mathcal{W}^{u}$. We first start by expressing the $cs$-conditionals $(\mu^{cs}_{J^{c}_{f},x})_{x\in\mathbb{T}^{3}}$ using the resonant state $\theta$. By (1.6), one has $\mathrm{WF}(\theta)\subset E_{cu}^{*}$ and for any $x\in\mathbb{T}^{3}$, $\mathrm{WF}([\mathcal{W}^{cs}(x)])\subset E_{cs}^{*}$, where $\mathcal{W}^{cs}(x)]$ denotes the integration current on $\mathcal{W}^{cs}(x)$, see for instance [undefab, Example 4.1.5]. Since $E_{cu}^{*}\cap E_{cs}^{*}=\emptyset$, this means that the distributional restriction $\theta_{x}$ where for any $g\in C^{\infty}(\mathbb{T}^{3})$, $\theta_{x}(g)=(\theta,g[\mathcal{W}^{cs}(x)])_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}$ is well defined [undefab, Corollary 4.2.2]. We first show that $(\theta_{x})_{x\in\mathbb{T}^{3}}$ defines a system of $cs$-measures.
Proposition 5.2.
The system $(\theta_{x})_{x\in\mathbb{T}^{3}}$ satisfies:
  • for any $x\in\mathbb{T}^{3}$, $\theta_{x}$ is a non-negative measure on $\mathcal{W}^{cs}(x)$;
  • for any $x\in\mathbb{T}^{3}$ and $y\in\mathcal{W}^{cs}(x)$, one has $\theta_{x}=\theta_{y}$;
  • for any $x\in\mathbb{T}^{3}$, one has
    (5.2) $f_{*}\theta_{x}=\lambda_{u}\theta_{f(x)};$
  • for any rectangle $R$ and $x,y\in R$, one has
    (5.3) $(\mathrm{Hol}^{u}_{x,y})_{*}\theta_{y}=\theta_{x}.$
We say that $(\theta_{x})_{x\in\mathbb{T}^{3}}$ is a center stable Margulis system of measures.
Proof.
Let $\omega\in\Omega_{2}$ be a smooth $2$-form which is $C^{0}$-close to $d\mathrm{Vol}_{\mathcal{W}^{cs}}$. This means that there is a function $a\in C^{0}(\mathbb{T}^{3})$, $C^{0}$-close to $1$ such that for any $g\in C^{\infty}(\mathbb{T}^{3})$,
$\forall n\in\mathbb{N},\quad(f_{*}^{n}\omega,g[\mathcal{W}^{cs}(x)])_{\Omega_{2}\times\mathcal{D}^{\prime}_{1}}=\int_{\mathcal{W}^{cs}(x)}g(y)a(f^{-n}y)d\big((f_{*})^{n}\mathrm{vol}_{\mathcal{W}^{cs}(x)}\big)(y).$
Since $\mathrm{WF}(g[\mathcal{W}^{cs}(x)])\subset E_{cs}^{*}\subset E_{s}^{*}$, one has $g[\mathcal{W}^{cs}(x)]\in\mathcal{H}^{-s}$ for some $s>0$, see [undefs, Lemma 1.11]. Lemma 4.7 then gives, for some $c>0$,
(5.4) $\begin{split}\theta_{x}(g)&=c\lim_{n\to+\infty}\lambda_{u}^{-n}(f_{*}^{n}\omega,g[\mathcal{W}^{cs}(x)])_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}\\ &=c\lim_{n\to+\infty}\lambda_{u}^{-n}\int_{\mathcal{W}^{cs}(x)}g(y)a(f^{-n}y)d\big((f_{*})^{n}\mathrm{vol}_{\mathcal{W}^{cs}(x)}\big)(y).\end{split}$
Since $a$ is $C^{0}$-close to $1$, one has $a>0$ and this shows that if $g\geq 0$, then $\theta_{x}(g)\geq 0$. In other words, $\theta_{x}$ is a non-negative measure on $\mathcal{W}^{cs}(x)$. The compatibility condition is immediate from the definition of $\theta_{x}$. Let us show (5.2). For this, we fix $g\in C^{\infty}(\mathbb{T}^{3})$. We use the relation $f^{*}(g[\mathcal{W}^{cs}(f(x))])=(g\circ f)[\mathcal{W}^{cs}(x)]$ to compute
$(f_{*}\theta_{x})(g)$ $=\theta_{x}(g\circ f)=(\theta,(g\circ f)[\mathcal{W}^{cs}(x)])_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}=(\theta,f^{*}(g[\mathcal{W}^{cs}(f(x))]))_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}$
$=(f_{*}\theta,g[\mathcal{W}^{cs}(f(x))])_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}=\lambda_{u}(\theta,g[\mathcal{W}^{cs}(f(x))])_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}=\lambda_{u}\theta_{f(x)}(g),$
where we used (1.6) for the resonance $\lambda_{u}$. We now show the invariance by unstable holonomy. The argument is given in [undefh, Section 4.1] but we include the proof here for completeness. Note that since $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, $f^{-1}$ is a partially hyperbolic diffeomorphism with dominated splitting $T(\mathbb{T}^{3})=E_{u}\oplus E_{cs}$ with $E_{u}$ being contracting for $f^{-1}$. We can thus use the absolute continuity of $E_{u}$, see [undefh, Theorem 4.3]. Namely, there are constants $K,\beta>0$ such that for any $\varepsilon>0$, for any rectangle $R$, for any $x_{1},x_{2}\in R$ and for any two $\varepsilon$-equivalent Borel sets $A_{i}\subset\mathcal{W}^{cs}(x_{i})$, $i=1,2$ (see §2.1.3), the measures $(\mathrm{Hol}^{u}_{x_{2},x_{1}})_{*}\mathrm{vol}_{A_{1}}$ and $\mathrm{vol}_{A_{2}}$ are equivalent and
(5.5) $\left|\frac{d(\mathrm{Hol}^{u}_{x_{2},x_{1}})_{*}\mathrm{vol}_{A_{1}}}{d\mathrm{vol}_{A_{2}}}-1\right|\leq K^{1+\varepsilon}\varepsilon^{\beta}.$
Let $R$ be a rectangle and $x,y\in R$. Let $\phi\in C^{\infty}(\mathcal{W}^{cs}(x))$ such that $\phi$ has compact support in $\mathcal{W}^{cs}(y)$ and let $\psi:=\phi\circ\mathrm{Hol}^{u}_{x,y}$. Suppose that $x,y$ are close enough so that $\psi$ and $\phi$ are $\delta$-equivalent for some small $\delta>0$. By the Anosov property, there is a $\nu>0$ such that for $n\in\mathbb{N}$, $\phi\circ f^{n}$ and $\psi\circ f^{n}$ are $(e^{-\nu n}\delta)$-equivalent via $\mathrm{Hol}^{u}_{f^{-n}x,f^{-n}y}=f^{-n}\circ\mathrm{Hol}^{u}_{x,y}\circ f^{n}$. In the following computation, we will write $\mathrm{vol}_{x}$ to denote $\mathrm{vol}_{\mathcal{W}^{cs}(x)}$. Using (5.5) and (3.4),
$\left|\int_{\mathcal{W}^{cs}(x)}\psi(z)a(f^{-n}z)d((f_{*})^{n}\mathrm{vol}_{x})(z)-\int_{\mathcal{W}^{cs}(y)}\phi(z)a(f^{-n}z)d((f_{*})^{n}\mathrm{vol}_{y})(z)\right|$
$=\left|\int_{\mathcal{W}^{cs}(f^{-n}x)}\psi(f^{n}z)a(z)d\mathrm{vol}_{f^{-n}x}(z)-\int_{\mathcal{W}^{cs}(f^{-n}y)}\phi(f^{n}z)a(z)d\mathrm{vol}_{f^{-n}y}(z)\right|$
$=\left|\int_{\mathcal{W}^{cs}(f^{-n}y)}\left(\phi\circ f^{n}\frac{d(\mathrm{Hol}^{u}_{f^{-n}y,f^{-n}x})_{*}\mathrm{vol}_{f^{-n}x}}{d\mathrm{vol}_{f^{-n}y}}-\phi\circ f^{n}\right)(z)a(z)d\mathrm{vol}_{f^{-n}y}(z)\right|$
$=\left|\int_{\mathcal{W}^{cs}(f^{-n}y)}\phi\circ f^{n}\left(\frac{d(\mathrm{Hol}^{u}_{f^{-n}y,f^{-n}x})_{*}\mathrm{vol}_{f^{-n}x}}{d\mathrm{vol}_{f^{-n}y}}-1\right)(z)a(z)d\mathrm{vol}_{f^{-n}y}(z)\right|$
$\leq K^{1+\delta e^{-\nu n}}(\delta e^{-\nu n})^{\beta}\int_{\mathcal{W}^{cs}(f^{-n}y)}|\phi\circ f^{n}(z)|a(z)d\mathrm{vol}_{f^{-n}y}(z)$
$\leq K^{\prime}\times K^{1+\delta e^{-\nu n}}(\delta e^{-\nu n})^{\beta}\lambda_{u}^{n}.$
Using (5.4), this shows that
[…]
which shows the holonomy invariance. ∎
Now, we show that $(\theta_{x})_{x\in\mathbb{T}^{3}}$ identifies with the center stable conditionals $(\mu^{cs}_{J^{c}_{f},x})_{x\in\mathbb{T}^{3}}$.
Proposition 5.3.
For any $x\in\mathbb{T}^{3}$,there is a $c_{x}>0$ such that for any $g\in C^{\infty}(\mathbb{T}^{3})$,
(5.6) $\lim_{n\to+\infty}((f^{*})^{n}\theta_{x})(g)=c_{x}\mu_{J^{c}_{f}}(g).$
Moreover, for any rectangle $R$ such that $\mu_{J^{c}_{f}}(R)>0$, the conditional measures $(\mu^{cs}_{J^{c}_{f},y})_{y\in R}$ are given $\mu_{J^{c}_{f}}$-a.e. by $(\theta_{y})_{y\in R}$, up to a constant rescaling. For $\mu_{J^{c}_{f}}$-a.e. $y\in R$, one has
(5.7) $\forall z\in\mathcal{W}^{cs}_{R}(y),\quad(\mu^{cs}_{J^{c}_{f},y})(z)=\frac{\theta_{y}(z)}{\theta_{y}(\mathcal{W}^{cs}_{R}(y))}.$
Proof.
Let $g\in C^{\infty}(\mathbb{T}^{3})$ and let $x\in\mathbb{T}^{3}$. We denote by $\Pi_{0}^{*}:\mathcal{H}^{-s}\to\mathcal{H}^{-s}$ the spectral projector on $\mathrm{Span}(\nu)$ where $\nu$ is a co-resonant state at $\lambda_{u}$. Using Lemma 4.7, we get $c_{x}\in\mathbb{R}$ such that $\lambda_{u}^{-n}(f^{*})^{n}[\mathcal{W}^{cs}(x)]\to c_{x}\nu$ in $\mathcal{H}^{-s}$ and Lemma 4.6 gives
$c_{x}\mu_{J^{c}_{f}}(g)$ $=c_{x}(g\theta,\nu)_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}=(g\theta,\Pi_{0}^{*}[\mathcal{W}^{cs}(x)])_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}$
$=\lim_{n\to+\infty}\lambda_{u}^{-n}(g\theta,(f^{*})^{n}[\mathcal{W}^{cs}(x)])_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}=\lim_{n\to+\infty}\lambda_{u}^{-n}((f_{*})^{n}(g\theta),[\mathcal{W}^{cs}(x)])_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}$
$=\lim_{n\to+\infty}\big(\big((f_{*})^{n}g\big)\theta,[\mathcal{W}^{cs}(x)]\big)_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}=\lim_{n\to+\infty}\big((f^{*})^{n}\theta_{x}\big)(g),$
where we used (1.6) for $\lambda=\lambda_{u}$ and $u=\theta$. The rest of the argument follows the proof of [undefk, Lemma 8.1]. We will need the following characterization of the conditional measures, see [undefk, Proposition 8.2].
Lemma 5.4.
Let $\mu$ be a probability measure and let $R\subset\mathbb{T}^{3}$ be a rectangle such that $\mu(R)>0$. Let $\{\xi_{n}\}_{n\in\mathbb{N}}$ be a sequence of refining finite partitions converging to the partition $\xi$ into center-stable manifolds $(\mathcal{W}^{cs}_{R}(x))_{x\in R}$. Then there is a $R^{\prime}\subset R$ such that $\mu(R)=\mu(R^{\prime})$ and such that for every $y\in R^{\prime}$ and every continuous $\psi:R\to\mathbb{R}$,
(5.8) $\int_{\mathcal{W}^{cs}_{R}(y)}\psi(z)d\mu^{cs}_{y}(z)=\lim_{n\to+\infty}\frac{1}{\mu(\xi_{n}(y))}\int_{\xi_{n}(y)}\psi(z)d\mu(z),$
where $\xi_{n}(y)$ is the element of $\xi_{n}$ containing $y$.
Let $x\in\mathbb{T}^{3}$, $n\geq 0$ and $R\subset\mathbb{T}^{3}$ be a rectangle so that $\mu_{J^{c}_{f}}(R)>0$. Let $\{\xi_{n}\}_{n\in\mathbb{N}}$ be a sequence of refining finite partitions so that for any $y\in R$, $\xi_{n}(y)$ is a rectangle and $\cap_{n\in\mathbb{N}}\xi_{n}(y)=\mathcal{W}^{cs}_{R}(y)$. Note that $(f^{*})^{n}\theta_{x}$ is supported on $\mathcal{W}^{cs}(f^{-n}x)$. Fix $y\in R^{\prime}$ given by Lemma 5.4 and let $n,m\in\mathbb{N}$. Then there exists points $z^{(1)}_{n,m},\ldots,z^{(s(m))}_{n,m}\in\mathcal{W}^{cs}(f^{-n}x)\cap\xi_{m}(y)$ such that $\mathcal{W}^{cs}(f^{-n}x)\cap\xi_{m}=\cup_{i=1}^{s(m)}\mathcal{W}^{cs}_{R}(z^{(i)}_{n,m})$, where the union is disjoint [^4] Since our system $(\theta_{x})_{x\in\mathbb{T}^{3}}$ is defined on the whole center stable manifold $\mathcal{W}^{cs}(x)$ rather than the local one, any $\mathcal{W}^{cs}_{R}(z^{(i)}_{n,m})$ is fully included in $\mathcal{W}^{cs}(f^{-n}x)$ in contrast to [undefk, Proposition 8.2]. . We want to write $((f^{*})^{n}\theta_{x})_{|\xi_{m}}$ as a linear combination of measures supported on these sets. Using (5.2) gives, for any continuous $\psi:\xi_{m}(y)\to\mathbb{R}$,
(5.9) $\begin{split}\int_{\xi_{m}(y)}\psi(z)d((f^{*})^{n}\theta_{x})(z)&=\sum_{i=1}^{s(m)}\int_{f^{n}\mathcal{W}^{cs}_{R}(z_{n,m}^{(i)})}\psi(f^{-n}z)d\theta_{x}(z)\\ &=\lambda_{u}^{-n}\sum_{i=1}^{s(m)}\int_{\mathcal{W}^{cs}_{R}(z_{n,m}^{(i)})}\psi(z)d\theta_{z_{n,m}^{(i)}}(z).\end{split}$
For any $i=1,\ldots,s(m)$, using the $u$-holonomy invariance (5.3), we obtain
$\int_{\mathcal{W}^{cs}_{R}(z_{n,m}^{(i)})}\psi(z)d\theta_{z_{n,m}^{(i)}}(z)=\int_{\mathcal{W}^{cs}_{R}(y)}\psi(z)d\theta_{y}(z).$
Plugging this back into (5.9), this gives
$\int_{\xi_{m}(y)}\psi(z)d((f^{*})^{n}\theta_{x})(z)=\lambda_{u}^{-n}s(m)\int_{\mathcal{W}^{cs}_{R}(y)}\psi(z)d\theta_{y}(z).$
Choosing $\psi\equiv 1$, this gives using (5.6)
$\lambda_{u}^{-n}s(m)=\frac{[(f^{*})^{n}\theta_{x}](\xi_{m}(y))}{\theta_{y}(\mathcal{W}^{cs}_{R}(y))}\to_{n\to+\infty}\frac{c_{x}\mu_{J^{c}_{f}}(\xi_{m}(f))}{\theta_{y}(\mathcal{W}^{cs}_{R}(y))}.$
Combining the last two equations finally yields
$\frac{1}{\mu_{J^{c}_{f}}(\xi_{m}(y))}\int_{\xi_{m}(y)}\psi d\mu_{J^{c}_{f}}\to_{n\to+\infty}\frac{1}{\theta_{y}(\mathcal{W}^{cs}_{R}(y))}\int_{\mathcal{W}^{cs}_{R}(y)}\psi d\theta_{y},$
which, combined with (5.8), gives (5.7). ∎
5.2. Holonomy invariance of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$.
In this subsection, we show Proposition 5.1.
Proof.
Let $R$ be a rectangle such that $\mu_{J^{c}_{f}}(R)>0$. By Proposition 5.3, the $cs$-conditionals $(\mu^{cs}_{J^{c}_{f},x})_{x\in\mathbb{T}^{3}}$ are given, up to a constant rescalling by $(\theta_{x})_{x\in\mathbb{T}^{3}}$ and are $\mathrm{Hol}^{u}$-invariant. Let $\theta$ be chosen so that $(\theta_{x})_{x\in\mathbb{T}^{3}}=(\mu^{cs}_{J^{c}_{f},x})_{x\in\mathbb{T}^{3}}$. For $q\in R$, recall that we defined a measure $\tilde{\mu}_{q}$ on $\mathcal{W}^{u}_{R}(q)$ by (2.5). The system of measures $(\tilde{\mu}_{q})_{q\in R}$ is easily seen to be $\mathrm{Hol}^{cs}$-invariant
(5.10) $\forall q\in R,\ \forall p\in\mathcal{W}^{cs}_{R}(q),\quad(\mathrm{Hol}^{cs}_{p,q})_{*}\tilde{\mu}_{q}=\tilde{\mu}_{p}.$
Using (2.6), for any $\phi\in C_{c}^{\infty}(R,\mathbb{R})$, one has
$\mu_{J^{c}_{f}}(\phi)=\int_{\mathcal{W}_{R}^{u}(q)}\int_{\mathcal{W}_{R}^{cs}(y)}\phi(z)d\theta_{y}(z)d\tilde{\mu}^{u}_{q}(y).$
We now use (5.3) and (5.10) to obtain
$\mu_{J^{c}_{f}}(\phi)$ $=\int_{\mathcal{W}^{u}_{R}(q)}\int_{\mathcal{W}^{cs}_{R}(y)}\phi(z)d\theta_{y}(z)d\tilde{\mu}^{u}_{q}(y)=\int_{\mathcal{W}^{u}_{R}(q)}\int_{\mathcal{W}_{R}^{cs}(q)}\phi([z,y])d\theta_{q}(z)d\tilde{\mu}^{u}_{q}(y)$
$=\int_{\mathcal{W}_{R}^{cs}(q)}\int_{\mathcal{W}_{R}^{u}(y)}\phi(z)d\tilde{\mu}^{u}_{y}(z)d\theta_{q}(y).$
From (2.4) and the last equation, we see that $\tilde{\mu}_{q}=(\mu^{u}_{J^{c}_{f}})_{q}=c_{R}\mu^{u}_{q}$ for some constant $c_{R}$ that might depend on the rectangle $R$ but not on $q\in R$. Then (5.10) shows that $(\mu^{u}_{q})_{q\in\mathbb{T}^{3}}$ is invariant by center stable holonomy and this concludes the proof. ∎
Note that we showed that $\mu_{J^{c}_{f}}$ has a center-stable/strong unstable local product structure. For any rectangle $R$ and $q\in R$, there is $c_{R}>0$ such that
(5.11) $\forall\phi\in L^{1}(R,\mu_{J^{c}_{f}}),\quad\mu_{J^{c}_{f}}(\phi)=c_{R}\int_{\mathcal{W}^{u}_{R}(q)}\int_{\mathcal{W}^{cs}_{R}(y)}\phi(z)d\mu^{cs}_{y}(z)d\mu^{u}_{q}(y).$
5.3. The $u$-Margulis system as a co-resonant state.
In this subsection, we show that the system of Margulis measures constructed in Theorem 3.4 coincides with the co-resonant state associated to the first resonance $\lambda_{u}$. Let $\varphi\in\Omega_{2}$ supported in a small rectangle $R$. We define a current by
(5.12) $\forall q\in R,\quad\mu_{q}^{u}(\varphi):=\int_{\mathcal{W}_{R}^{u}(q)}\Big(\int_{\mathcal{W}^{cs}_{R}(y)}\varphi\Big)d\mu^{u}_{q}(y).$
Lemma 5.5.
The current $\mu^{u}_{q}$ is independent of $q\in R$ and thus defines a current on $\mathbb{T}^{3}$ which we will denote by $\mu^{u}$. Moreover, one has
(5.13) $f^{*}\mu^{u}=\lambda_{u}\mu^{u},\quad\mathrm{WF}(\mu^{u})\subset E_{cs}^{*}\subset E_{s}^{*}.$
In other words, $\mu^{u}$ is a co-resonant state associated to the first resonance $\lambda_{u}$.
Proof.
Let $p,q\in R$. If $p\in\mathcal{W}^{u}(q)$, then $\mu^{u}_{p}=\mu^{u}_{q}$ which implies that $\mu^{u}_{p}(\varphi)=\mu^{u}_{q}(\varphi)$. If $q\in\mathcal{W}^{cs}(p)$, then using the holonomy invariance (5.1), we have
$\mu_{q}^{u}(\varphi)=\int_{\mathcal{W}^{u}_{R}(q)}\Big(\int_{\mathcal{W}^{cs}_{R}(y)}\varphi\Big)d\mu^{u}_{q}(y)=\int_{\mathcal{W}_{R}^{u}(p)}\Big(\int_{\mathcal{W}^{cs}_{R}(z)}\varphi\Big)d\big((\mathrm{Hol}^{cs}_{p,q})_{*}\mu^{u}_{q}\big)(z)=\mu_{p}^{u}(\varphi).$
Using the local product structure, we deduce that $\mu^{u}_{q}$ does not depend on the point $q\in R$. Using a partition of unity $(\chi_{i})_{i=1}^{m}$ associated to a finite cover $(R_{i})_{i=1}^{m}$ by rectangles, we can define $\mu^{u}$ globally on $\mathbb{T}^{3}$:
$\forall\varphi\in\Omega_{2},\quad\mu^{u}(\varphi)=\sum_{i=1}^{m}\mu^{u}_{p_{i}}(\chi_{i}\varphi),\ p_{i}\in R,$
and the value does not depend on any choice. The first equation in (5.13) is a consequence of (3.8). Indeed, let $\varphi\in\Omega_{2}$ be such that $\varphi$ is supported in $R_{i}$ and $f_{*}\varphi$ is supported in $R_{j}$, then for some $p_{i}\in R_{i}$ and $p_{j}\in R_{j}$,
$f^{*}\mu^{u}(\varphi)$ $=\mu^{u}(f_{*}\varphi)=\int_{\mathcal{W}_{R_{j}}^{u}(p_{j})}\Big(\int_{\mathcal{W}^{cs}_{R_{j}}(y)}f_{*}\varphi\Big)d\mu^{u}_{p_{j}}(y)=\int_{\mathcal{W}_{R_{j}}^{u}(p_{j})}\Big(\int_{f^{-1}(\mathcal{W}^{cs}_{R_{j}}(y))}\varphi\Big)d\mu^{u}_{p_{j}}(y)$
$=\int_{\mathcal{W}_{R_{j}}^{u}(p_{j})}\Big(\int_{\mathcal{W}_{R_{j}}^{cs}(f^{-1}(y))}\varphi\Big)d\mu^{u}_{p_{j}}(y)=\int_{\mathcal{W}_{R_{i}}^{u}(p_{i})}\Big(\int_{\mathcal{W}_{R_{i}}^{cs}(z)}\varphi\Big)d(f^{*}\mu^{u}_{p_{j}})(y)$
$=\lambda_{u}\int_{\mathcal{W}_{R_{i}}^{u}(p_{i})}\Big(\int_{\mathcal{W}_{R_{i}}^{cs}(z)}\varphi\Big)d\mu^{u}_{p_{i}}(y)=\lambda_{u}\mu^{u}(\varphi).$
We are left with showing the wavefront set bound. The proof follows [undefz, Lemma 3.2]. Let $\chi$ be a smooth $2$-form supported in $R$ and $S\in C^{\infty}$ such that $dS(q)=\xi\notin E_{cs}^{*}.$
$\mu^{u}(e^{i\frac{S}{h}}\chi)=\int_{\mathcal{W}^{u}_{R}(q)}\Big(\int_{\mathcal{W}_{R}^{cs}(x)}e^{i\frac{S(y)}{h}}\chi(y)\Big)d\mu^{u}_{q}(x).$
Integrating by parts in $y$ shows that the integrand is $O(h^{\infty})$ as long as $dS$ does not vanish on $\mathcal{W}_{R}^{cs}(x)$, which can be ensured near $q$ by the definition of $E_{cs}^{*}$, see (4.5). This shows that $\xi\notin\mathrm{WF}(\mu^{u})$ and thus $\mathrm{WF}(\mu^{u})\subset E_{cs}^{*}$. ∎
Remark 3.
As explained in Remark 1, the condition $\mathrm{WF}({\mu^{u}})\subset E_{cs}^{*}$ is stronger than the condition in (1.7). Lemma 4.6 and [undefab, Lemma 4.3.2] imply that that $\mathrm{WF}(\mu_{J^{c}_{f}})\subset E_{cs}^{*}\oplus E_{cu}^{*}$. This is better than the usual wavefront set bound $\mathrm{WF}(\mu)\subset E_{s}^{*}\oplus E_{cu}^{*}$ for an equilibrium state $\mu$. This extra regularity in the center direction is actually a consequence of the fact that the conditional densities of $\mu_{J^{c}_{f}}$ along the center direction are absolutely continuous with respect to the Lebesgue measure. This should be compared to the case of the SRB measure for which $\mathrm{WF}(\mu_{\mathrm{SRB}})\subset E_{cu}^{*}$ since its conditional densities on the stable manifolds are absolutely continuous.
Proposition 5.6.
The conditionals of $\mu_{J^{c}_{f}}$ on the center foliation $(\mathcal{W}^{c}(x))_{x\in\mathbb{T}^{3}}$ are absolutely continuous with respect to the Lebesgue measures $d\mathrm{Vol}_{\mathcal{W}^{c}(x)}$.
Proof.
For $z\in\mathbb{T}^{3}$ and $y\in\mathcal{W}^{c}(z)$, we define
(5.14) $\omega(z,y):=\sum_{k=0}^{+\infty}\big(J^{c}_{f}(f^{-k}y)-J^{c}_{f}(f^{-k}z)\big),$
where the sum converges since $\mathcal{W}^{c}$ is uniformly expanding and $J^{c}_{f}$ is Hölder continuous. We first use that the strong unstable foliation is $C^{1}$ when restricted to the unstable foliation, see for instance [undef, Section 2.2.1]. In particular, that there is $\delta>0$, such that for any $x\in\mathbb{T}^{3}$ and $y\in\mathcal{W}^{u}(x,\delta)$, one has $(\mathrm{Hol}^{u}_{x,y})_{*}\mathrm{Vol}_{\mathcal{W}^{c}(x)}\ll\mathrm{Vol}_{\mathcal{W}^{c}(y)}$. The density can be computed using the fact that $d(f^{-k}x,f^{-k}y)\to 0$ when $k\to+\infty$ and the fact that the holonomy is $C^{1}$. A direct computation using Jacobians shows that for any $z\in\mathbb{T}^{3}$ and $y\in\mathcal{W}^{c}(z)$, the density is given by the exponential of the function $\omega$ defined in (5.14)
(5.15) $\forall u\in\mathcal{W}^{c}(y),\quad\frac{d\big((\mathrm{Hol}^{u}_{z,y})_{*}\mathrm{Vol}_{\mathcal{W}^{c}(z)}\big)}{d\mathrm{Vol}_{\mathcal{W}^{c}(y)}}(u)=e^{\omega(u,\mathrm{Hol}^{s}_{y,z}u)}.$
We now define a system of measures on the unstable foliation $\mathcal{W}^{cu}$. For any $x\in\mathbb{T}^{3}$ and any compactly supported $\phi\in C(\mathcal{W}^{cu}(x))$, we let
(5.16) $\mu^{cu}_{x}(\phi)=\int_{\mathcal{W}^{u}(x)}\left(\int_{\mathcal{W}^{c}(y)}e^{\omega(z,y)}\phi(z)d\mathrm{Vol}_{\mathcal{W}^{c}(y)}(z)\right)d\mu^{u}_{x}(y).$
The definition does not depend on the choice of base point $x$ in the unstable leaf by (5.15) and the fact that $\mu^{u}$ does not depend on the choice of base point in the strong unstable leaf. Since $\mu^{u}$ was shown to be a Radon measure with full support on each strong unstable leaf, see Theorem 2, this shows that $\mu^{cu}$ is also a Radon measure with full support on each unstable leaf. We now compute, using $J^{c}_{f}(z)=\omega(f(z),z)$ and the cocycle relation $\omega(x,y)+\omega(y,z)=\omega(x,z)$,
$\mu^{cu}_{x}(e^{-J^{c}_{f}(\cdot)}\phi\circ f)$ $=\int_{\mathcal{W}^{u}(x)}\left(\int_{\mathcal{W}^{c}(y)}e^{\omega(z,y)}e^{-J^{c}_{f}(z)}\phi(f(z))d\mathrm{Vol}_{\mathcal{W}^{c}(y)}(z)\right)d\mu^{u}_{x}(y)$
$=\int_{\mathcal{W}^{u}(x)}\left(\int_{\mathcal{W}^{c}(y)}e^{\omega(z,y)}e^{\omega(f(z),z)}\phi(f(z))d\mathrm{Vol}_{\mathcal{W}^{c}(y)}(z)\right)d\mu^{u}_{x}(y)$
$=\int_{\mathcal{W}^{u}(x)}\left(\int_{\mathcal{W}^{c}(y)}e^{\omega(f(z),y)}\phi(f(z))d\mathrm{Vol}_{\mathcal{W}^{c}(y)}(z)\right)d\mu^{u}_{x}(y)$
$=\int_{\mathcal{W}^{u}(x)}f^{*}\big(\mathrm{Vol}_{\mathcal{W}^{c}(y)}(e^{\omega(\cdot,y)}\phi(\cdot))\big)d\mu^{u}_{x}(y)$
$=\int_{\mathcal{W}^{u}(f(x))}\left(\int_{\mathcal{W}^{c}(y)}e^{\omega(z,y)}e^{-J^{c}_{f}(z)}\phi(f(z))d\mathrm{Vol}_{\mathcal{W}^{c}(y)}(z)\right)df_{*}(\mu^{u}_{x})(y)$
$=e^{-h^{u}_{\mathrm{top}}(f)}\mu_{f(x)}^{cu}(\phi)=e^{-P(J^{c}_{f})}\mu_{f(x)}^{cu}(\phi),$
where we used that $h^{u}_{\mathrm{top}}(f)=P(J^{c}_{f})$ by Theorem 1 and the conformal property of $\mu^{u}$, see Theorem 2.
In other words, we checked that the system $(\mu^{cu}_{x})_{x\in\mathbb{T}^{3}}$ is a $cu$-system which satisfies properties $(2)$ and $(3)$ of [undefi, Theorem A] for the potential $J^{c}_{f}$. We can thus apply [undefi, Theorem B] which gives that the $cu$-conditionals of the equilibrium state $\mu_{J^{c}_{f}}$ associated to $J^{c}_{f}$ are absolutely continuous with respect to $\mu^{cu}$.
From (5.15), the $c$-conditionals of $\mu^{cu}$ are absolutely continuous with respect to the Lebesgue measure on $\mathcal{W}^{c}$ and we deduce that this is also the case for $\mu_{J^{c}_{f}}$. ∎